### v1

# Determination of the stably free cancellation property for orders

Abstract.
Let $K$ be a number field, let $A$ be a finite-dimensional semisimple $K$-algebra, and let $\Lambda$ be an $\mathcal{O}_{K}$-order in $A$. We give practical algorithms that determine whether or not $\Lambda$ has the stably free cancellation property (SFC). As an application, we determine all finite groups $G$ of order at most $383$ such that the integral group ring $\mathbb{Z}[G]$ has SFC.
Key words and phrases:
stably free cancellation, locally free cancellation
2020 Mathematics Subject Classification:
16H10, 16H20, 16Z05, 20C05, 20C10, 11R33, 11Y40
1. Introduction
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$. Let $A$ be a finite-dimensional semisimple $K$-algebra and let $\Lambda$ be an $\mathcal{O}$-order in $A$. For example, if $G$ is a finite group then the group ring $\mathcal{O}[G]$ is an order in the group algebra $K[G]$. For $k\geq 0$, let $\Lambda^{(k)}$ denote the (left) $\Lambda$-module that is the direct sum of $k$ copies of $\Lambda$. The order $\Lambda$ is said to have stably free cancellation (SFC) if every finitely generated stably free $\Lambda$-module is in fact free, that is, if $X$ is a finitely generated $\Lambda$-module such that $X\oplus\Lambda^{(n)}\cong\Lambda^{(n+m)}$ for some $m,n\geq 0$, then $X\cong\Lambda^{(m)}$.
One can also consider locally free cancellation (LFC), the definition of which we defer to §3.1. If $\Lambda$ has LFC then it also has SFC. Note, however, that there exist orders that have SFC but do not have LFC (see [Sme15] or [SV19]). We say that $A$ satisfies the Eichler condition if no simple component of $A$ is a totally definite quaternion algebra (see §3.2). Jacobinski [Jac68] showed that if $A$ satisfies the Eichler condition then $\Lambda$ has LFC and thus SFC. Smertnig–Voight [SV19] showed that in the case that $A$ is a totally definite quaternion algebra, the question of whether $\Lambda$ has SFC or LFC can be decided algorithmically. They then used this to classify all orders in totally definite quaternion algebras that have SFC, and of these, all those that have LFC. This classification builds on work of Vignéras [Vig76], Hallouin–Maire [HM06] and Smertnig [Sme15].
For notational convenience, we henceforth abuse language by referring to direct sums of orders and algebras rather than direct products. We can always decompose $A$ into a direct sum of $K$-algebras $A=A_{1}\oplus A_{2}\oplus\cdots\oplus A_{k}$, where $A_{1}$ satisfies the Eichler condition and $A_{2},\ldots,A_{k}$ are all totally definite quaternion algebras. If $\Lambda$ decomposes as $\Lambda=\Lambda_{1}\oplus\Lambda_{2}\oplus\cdots\oplus\Lambda_{k}$ such that each $\Lambda_{i}$ is contained in $A_{i}$, then the question of whether $\Lambda$ has SFC or LFC can be decided using the results discussed in the previous paragraph, since both properties respect direct sums. However, in many cases of interest, $\Lambda$ does not exhibit such a decomposition.
In this article, we show that the question of whether an arbitrary order $\Lambda$ has SFC or not can in principle always be decided algorithmically. In fact, we present three algorithms, all of which are practical in certain situations. The first of these, Algorithm 8.9, is only practical when $A$ is of ‘small’ dimension. The second, Algorithm 9.1, cannot prove that $\Lambda$ has SFC, but can be used to show that it fails SFC, and in this situation is often much faster than Algorithm 8.9. The third, Algorithm 10.3, uses fibre products to reduce the problem for the original order to that for an order in an algebra of smaller dimension as well as the computation of a certain subgroup of the unit group of a finite ring. These algorithms require several other algorithms of independent interest, including those that can be used to compute certain primary decompositions and unit groups of finite rings (see §11), and improved freeness tests for certain modules over orders (see §12).
Orders of particular interest are integral group rings $\mathbb{Z}[G]$ for $G$ a finite group. For instance, the property of whether or not $\mathbb{Z}[G]$ has SFC or LFC has applications in topology (see [Joh03], [Nic21b], [Nic20], for example) and in the determination of whether a finite Galois extension of $\mathbb{Q}$ has a normal integral basis or not (see [Cou94], [Cou98]).
Let $G$ be a finite group and let $\mathbb{H}$ denote the quaternion division algebra over $\mathbb{R}$. If no simple component of $\mathbb{R}[G]$ is isomorphic to $\mathbb{H}$, then $\mathbb{Q}[G]$ satisfies the Eichler condition and so $\mathbb{Z}[G]$ has LFC (and thus SFC) by the aforementioned result of Jacobinski. This condition on $G$ is equivalent to the requirement that $G$ has no quotient that is isomorphic to a binary polyhedral group, that is, a generalised quaternion group $Q_{4n}$ for $n\geq 2$, or one of $\tilde{T}$, $\tilde{O}$ or $\tilde{I}$, the binary tetrahedral, octahedral and icosahedral groups. In §4, we review important work of Swan [Swa83], Chen [Che86], and Nicholson [Nic21a] on SFC and LFC for $\mathbb{Z}[G]$ in the case that $G$ does have a quotient isomorphic to a binary polyhedral group.
Before stating our own results on SFC for integral group rings, we first establish some further notation. Let $C_{n}$ denote the cyclic group of order $n$. Let $G_{(n,m)}$ denote the $m$th group of order $n$ in the GAP [GAP24] small groups library [BEO02].
A special case of a result due to Swan and Fröhlich says that if $\mathbb{Z}[G]$ has SFC then $\mathbb{Z}[G/N]$ has SFC for every quotient $G/N$, or equivalently, if $\mathbb{Z}[G/N]$ fails SFC for some choice of $N$, then $\mathbb{Z}[G]$ also fails SFC (see Theorem 3.1 and Lemma 4.1). The following result is a special case of Theorem 9.4, in which the claim that each $\mathbb{Z}[G]$ fails SFC was proven using Algorithm 9.1; the claim that $\mathbb{Z}[G/N]$ has SFC for each proper quotient $G/N$ can be interpreted as saying that each $G$ is minimal with respect to quotients.
Theorem 1.1.
Let $G$ be one of the following finite groups:
  1. (i)
    $Q_{4n}\times C_{2}$ for $2\leq n\leq 5$,
  2. (ii)
    $\tilde{T}\times C_{2}^{2}$, $\tilde{O}\times C_{2}$, $\tilde{I}\times C_{2}$, or
  3. (iii)
    $G_{(32,14)}$, $G_{(36,7)}$, $G_{(64,14)}$, $G_{(100,7)}$.
Then $\mathbb{Z}[G]$ fails SFC but $\mathbb{Z}[G/N]$ has SFC for every proper quotient $G/N$.
Note that Swan [Swa83] already proved that $\mathbb{Z}[Q_{8}\times C_{2}]$ fails SFC and Chen [Che86] already proved that $\mathbb{Z}[G]$ fails SFC for $G=Q_{12}\times C_{2}$, $Q_{16}\times C_{2}$, $Q_{20}\times C_{2}$, $G_{(36,7)}$, $G_{(100,7)}$. Since Swan [Swa83] proved that $\mathbb{Z}[Q_{4n}]$ fails SFC for $n\geq 6$, we obtain the following, which is the most general result to date on integral group rings that fail SFC.
Corollary 1.2.
Let $G$ be a finite group with a quotient isomorphic to a group listed in Theorem 1.1 or to $Q_{4n}$ for some $n\geq 6$. Then $\mathbb{Z}[G]$ fails SFC.
Using Algorithm 10.3 with judicious choices of fibre products, we prove the following result on new cases of integral group rings with SFC (this is Theorem 13.1).
Theorem 1.3.
Let $G$ be one of the following finite groups:
$\tilde{T}\times C_{2},\,\tilde{T}\times Q_{12},\,\tilde{T}\times Q_{20},\,Q_{8}\rtimes\tilde{T},\,G_{(96,66)},\,G_{(192,183)},\,G_{(480,962)}.$
Then $\mathbb{Z}[G]$ has SFC.
By combining Corollary 1.2 and Theorem 1.3, and using the GAP [GAP24] small groups library [BEO02], we obtain the following result (see Corollary 13.3).
Theorem 1.4.
Let $G$ be a finite group with $|G|\leq 383$. Then $\mathbb{Z}[G]$ has SFC if and only if $G$ has no quotient of the form
  1. (i)
    $Q_{4n}\times C_{2}$ for $2\leq n\leq 5$,
  2. (ii)
    $\tilde{T}\times C_{2}^{2}$, $\tilde{O}\times C_{2}$, $\tilde{I}\times C_{2}$,
  3. (iii)
    $G_{(32,14)}$, $G_{(36,7)}$, $G_{(64,14)}$, $G_{(100,7)}$, or
  4. (iv)
    $Q_{4n}$ for $6\leq n\leq 95$.
This represents a significant advance in the state-of-the-art, since before the results of this article, such a classification was only known for groups $G$ with $|G|\leq 31$ (for example, it was not known whether $\mathbb{Z}[G_{(32,14)}]$ had SFC or not). In fact, Theorem 1.4 is a special case of Theorem 13.2, which gives a classification for $|G|\leq 1023$, but with $27$ explicitly listed gaps. In other words, of the $11759892$ groups $G$ with $|G|\leq 1023$, there are now only $27$ for which it is not known whether $\mathbb{Z}[G]$ has SFC or not.
Nicholson [Nic24, Theorem 1.7] has now shown that $\mathbb{Z}[\tilde{T}\times C_{2}]$ in fact has LFC, using our result that it has SFC (i.e. part of Theorem 1.3) as an input. By combining this with further new results on integral group rings with LFC, and using Theorem 1.1 as a crucial input, he has given classifications of all integral group rings $\mathbb{Z}[G]$ with LFC and SFC, provided that either (i) $G$ has no quotient isomorphic to $\tilde{O}$ or $\tilde{I}$, or (ii) $G$ does have a quotient isomorphic to $C_{2}\times C_{2}$ (see [Nic24, Theorems A and B]). We remark that in the case of integral group rings, the LFC property discussed here is equivalent to the projective cancellation (PC) property discussed in loc. cit.
Acknowledgements
The authors are grateful to John Nicholson for helpful comments and discussions, including suggestions for which integral group rings to investigate initially. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any author accepted manuscript version arising.
Funding
Hofmann gratefully acknowledges support by the Deutsche Forschungsgemeinschaft — Project-ID 286237555 – TRR 195; and Project-ID 539387714.
2. Preliminaries on lattices and orders
For further background on lattices and orders, we refer the reader to [CR81, Rei03]. Henceforth, all rings considered will be associative and unital.
2.1. Lattices and orders
Let $R$ be a Dedekind domain with field of fractions $K$. An $R$-lattice is a finitely generated torsion-free $R$-module, or equivalently, a finitely generated projective $R$-module. For any finite-dimensional $K$-vector space $V$, an $R$-lattice in $V$ is a finitely generated $R$-submodule $M$ in $V$. We define a $K$-vector subspace of $V$ by
$KM:=\{\alpha_{1}m_{1}+\alpha_{2}m_{2}+\cdots+\alpha_{r}m_{r}\mid r\in\mathbb{Z}_{\geq 0},\alpha_{i}\in K,m_{i}\in M\}$
and say that $M$ is a full $R$-lattice in $V$ if $KM=V$. We may identify $KM$ with $K\otimes_{R}M$.
Now let $A$ be a finite-dimensional semisimple $K$-algebra and let $\Lambda$ be an $R$-order in $A$, that is, a subring of a $A$ that is an $R$-lattice such that $K\Lambda=A$. A $\Lambda$-lattice is a left $\Lambda$-module that is also an $R$-lattice.
2.2. Localisations and completions
Henceforth assume that $K$ is a number field.
Let $\mathfrak{p}$ range over all maximal ideals of $R$. Let $R_{\mathfrak{p}}$ denote the localisation of $R$ at $\mathfrak{p}$, that is, the ring of $\mathfrak{p}$-integral elements in $K$. For an $R$-lattice $M$, define the localisation $M_{\mathfrak{p}}$ to be the $R_{\mathfrak{p}}$-lattice $R_{\mathfrak{p}}\otimes_{R}M$. Localising a $\Lambda$-lattice $X$ at $\mathfrak{p}$ yields a $\Lambda_{\mathfrak{p}}$-lattice $X_{\mathfrak{p}}$. Two $\Lambda$-lattices $X$ and $Y$ are said to be locally isomorphic, or in the same genus, if there is an isomorphism of $\Lambda_{\mathfrak{p}}$-lattices $X_{\mathfrak{p}}\cong Y_{\mathfrak{p}}$ for every $\mathfrak{p}$, in which case we write $X\vee Y$.
Let $\widehat{R}_{\mathfrak{p}}$ denote the $\mathfrak{p}$-adic completion of $R$. For an $R$-lattice $M$, define the completion $\widehat{M}_{\mathfrak{p}}$ to be the $\widehat{R}_{\mathfrak{p}}$-lattice $\widehat{R}_{\mathfrak{p}}\otimes_{R}M$. Completing a $\Lambda$-lattice $X$ at $\mathfrak{p}$ yields a $\widehat{\Lambda}_{\mathfrak{p}}$-lattice $\widehat{X}_{\mathfrak{p}}$. Given $\Lambda$-lattices $X$ and $Y$, we have that $X_{\mathfrak{p}}\cong Y_{\mathfrak{p}}$ as $\Lambda_{\mathfrak{p}}$-lattices if and only if $\widehat{X}_{\mathfrak{p}}\cong\widehat{Y}_{\mathfrak{p}}$ as $\widehat{\Lambda}_{\mathfrak{p}}$-lattices (see [CR81, (30.17)] or [Rei03, (18.2)]). Therefore $X\vee Y$ if and only if there is an isomorphism of $\widehat{\Lambda}_{\mathfrak{p}}$-lattices $\widehat{X}_{\mathfrak{p}}\cong\widehat{Y}_{\mathfrak{p}}$ for every $\mathfrak{p}$.
2.3. Locally free lattices and the principal genus
A $\Lambda$-lattice $X$ is said to be locally free of rank $n$ if $X\vee\Lambda^{(n)}$, where $\Lambda^{(n)}$ denotes the direct sum of $n$ copies of $\Lambda$. Note that every locally free $\Lambda$-lattice is projective over $\Lambda$ by [CR81, (8.19)].
Define the principal genus $g(\Lambda)$ to be the collection of locally free $\Lambda$-lattices of rank $1$. By [CR81, (31.7)], given $Y_{1},Y_{2},Y_{3}\in g(\Lambda)$, there exists $Y_{4}\in g(\Lambda)$ such that
$Y_{1}\oplus Y_{2}\cong Y_{3}\oplus Y_{4}.$ (2.1)
Moreover, by [CR81, (31.14)], given $Y\vee\Lambda^{(n)}$, there exists $X\in g(\Lambda)$ such that
$Y\cong\Lambda^{(n-1)}\oplus X.$ (2.2)
For $X\in g(\Lambda)$ we write $[X]_{\mathrm{iso}}$ for the isomorphism class of $X$ and define
$\mathrm{LF}_{1}(\Lambda)=\{[X]_{\mathrm{iso}}:X\in g(\Lambda)\}.$
Note that $\mathrm{LF}_{1}(\Lambda)$ is finite by virtue of the Jordan–Zassenhaus theorem [Rei03, (26.4)].
2.4. Stable isomorphism
Let $P(\Lambda)$ denote the collection of finitely generated projective $\Lambda$-modules. For $X,Y\in P(\Lambda)$, we say that $X$ and $Y$ are stably isomorphic if
$X\oplus\Lambda^{(m)}\cong Y\oplus\Lambda^{(m)}\textrm{ for some }m\geq 0.$
A module $X\in P(\Lambda)$ is said to be stably free of rank $n$ if it is stably isomorphic to $\Lambda^{(n)}$. For $X\in P(\Lambda)$, let $[X]_{\mathrm{st}}$ denote the stable isomorphism class of $X$. Note that by [CR81, (6.15)] and the discussion in §2.2, stably isomorphic $\Lambda$-lattices are locally isomorphic. In particular, stably free lattices are locally free.
2.5. Locally free class groups
We give an overview of the locally free class group of $\Lambda$, for which there are several equivalent definitions (see [CR87, §49A]). Here we follow [CR87, (49.10)], to which we refer the reader for further details. Let
$\mathrm{Cl}(\Lambda)=\{[X]_{\mathrm{st}}:X\in g(\Lambda)\}.$
By (2.1), given $X_{1},X_{2}\in g(\Lambda)$ there exists $X_{3}\in g(\Lambda)$ such that $X_{1}\oplus X_{2}\cong\Lambda\oplus X_{3}$. It is straightforward to check that the stable isomorphism class $[X_{3}]_{\mathrm{st}}$ is uniquely determined by the classes $[X_{1}]_{\mathrm{st}}$ and $[X_{2}]_{\mathrm{st}}$. We define a binary operation on $\mathrm{Cl}(\Lambda)$ by
$[X_{1}]_{\mathrm{st}}+[X_{2}]_{\mathrm{st}}=[X_{3}]_{\mathrm{st}}\textrm{ whenever }X_{1}\oplus X_{2}\cong\Lambda\oplus X_{3},$
where $X_{1},X_{2},X_{3}\in g(\Lambda)$. This is now a well-defined commutative and associative operation, with identity element $[\Lambda]_{\mathrm{st}}$. Moreover, inverses exist since by (2.1), given $X\in g(\Lambda)$ there exists $X^{\prime}\in g(\Lambda)$ such that $X\oplus X^{\prime}\cong\Lambda^{(2)}$. Therefore $\mathrm{Cl}(\Lambda)$ is an abelian group, and is finite since $\mathrm{LF}_{1}(\Lambda)$ is finite.
3. Review of cancellation properties for orders
Let $R$ be a Dedekind domain whose quotient field $K$ is a number field. Henceforth every $R$-order is assumed to be in a finite-dimensional semisimple $K$-algebra.
3.1. Stably free and locally free cancellation
Let $\Lambda$ be an $R$-order. Then $\Lambda$ is said to have stably free cancellation (SFC) if every finitely generated stably free $\Lambda$-module is in fact free, that is, if for every finitely generated $\Lambda$-module $X$ we have
$X\oplus\Lambda^{(m)}\cong\Lambda^{(m+n)}\textrm{ for some }m,n\geq 0\implies X\cong\Lambda^{(n)}.$ (3.1)
Similarly, $\Lambda$ is said to have locally free cancellation (LFC) if for every pair of locally free $\Lambda$-lattices $X$ and $Y$ we have
$X\oplus\Lambda^{(m)}\cong Y\oplus\Lambda^{(m)}\textrm{ for some }m\geq 0\implies X\cong Y.$ (3.2)
Since locally free $\Lambda$-lattices are projective (see §2.3), one can easily deduce that $\Lambda$ has LFC if and only if it has the apparently stronger property that justifies the name: for every triple of locally free $\Lambda$-lattices $X,Y,Z$ we have
$X\oplus Z\cong Y\oplus Z\implies X\cong Y.$
If $X$ is a finitely generated $\Lambda$-module such that $X\oplus\Lambda^{(m)}\cong\Lambda^{(m+n)}\textrm{ for some }m,n\geq 0$ then it is straightforward to see that $X$ is a $\Lambda$-lattice; since $X$ is stably free it is also locally free (see §2.4). Therefore if $\Lambda$ has LFC then it also has SFC. Note, however, that there exist orders that have SFC but do not have LFC (see [Sme15] or [SV19]).
By (2.2), $\Lambda$ has SFC if and only if (3.1) holds for all $X\in g(\Lambda)$. Similarly, $\Lambda$ has LFC if and only if (3.2) holds for all $X,Y\in g(\Lambda)$. Hence the canonical map $\mathrm{LF}_{1}(\Lambda)\rightarrow\mathrm{Cl}(\Lambda)$ defined by $[X]_{\mathrm{iso}}\mapsto[X]_{\mathrm{st}}$ is surjective by definition of $\mathrm{Cl}(\Lambda)$ and is bijective precisely when $\Lambda$ has LFC. Moreover, $\Lambda$ has SFC precisely when the fibre above 00 has exactly one element, that is, when every stably free $\Lambda$-lattice of rank $1$ is in fact free (of rank $1$).
Theorem 3.1 (Fröhlich, Swan).
Let $\Lambda\rightarrow\Gamma$ be a map of $R$-orders inducing a surjection $K\Lambda\rightarrow K\Gamma$ of $K$-algebras. If $\Lambda$ has SFC (resp. LFC) then $\Gamma$ has SFC (resp. LFC).
Proof.
The claim for LFC was proven by Fröhlich [Frö75, VIII], whose approach was subsequently refined by Swan. Here we adapt the proof of [Nic21a, Theorem 1.1]. By [Swa83, Theorem A10] the diagram
${\mathrm{LF}_{1}(\Lambda)}$${\mathrm{LF}_{1}(\Gamma)}$${\mathrm{Cl}(\Lambda)}$${\mathrm{Cl}(\Gamma)}$
is a weak pull back in the sense that a compatible pair of elements in $\mathrm{Cl}(\Lambda)$ and $\mathrm{LF}_{1}(\Gamma)$ has a not-necessarily-unique lift to $\mathrm{LF}_{1}(\Lambda)$. It follows that the fibre in $\mathrm{LF}_{1}(\Lambda)$ over an element $x\in\mathrm{Cl}(\Lambda)$ maps onto the fibre in $\mathrm{LF}_{1}(\Gamma)$ over the image of $x$ in $\mathrm{Cl}(\Gamma)$. Thus if $\Lambda$ has LFC then $\Gamma$ has LFC. Taking $x=0$ shows that if $\Lambda$ has SFC then $\Gamma$ has SFC. ∎
Corollary 3.2.
Let $\Lambda\subset\Gamma$ be $R$-orders such that $K\Lambda=K\Gamma$. If $\Lambda$ has SFC (resp. LFC) then $\Gamma$ has SFC (resp. LFC).
Corollary 3.3.
Suppose that $\Lambda=\Lambda_{1}\oplus\Lambda_{2}$ is a direct sum of two $R$-orders $\Lambda_{1}$ and $\Lambda_{2}$. Then $\Lambda$ has SFC (resp. LFC) if and only if both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC (resp. LFC).
Proof.
One direction follows from Theorem 3.1; the other follows from the definitions. ∎
3.2. Totally definite quaternion algebras and the Eichler condition
Let $\mathbb{H}$ denote the quaternion division algebra over $\mathbb{R}$. A totally definite quaternion algebra is a simple finite-dimensional $K$-algebra $B$ whose completion at every archimedean prime of the centre of $B$ is isomorphic to $\mathbb{H}$ (in particular, this forces $K$ to be totally real).
We say that a finite-dimensional semisimple $K$-algebra $A$ satisfies the Eichler condition if no simple component of $A$ is a totally definite quaternion algebra. A proof of the following result can be found in [CR87, (51.24)] or [Jac68, §4].
Theorem 3.4 (Jacobinski).
If $A$ is a finite-dimensional semisimple $K$-algebra satisfying the Eichler condition then every $R$-order in $A$ has LFC.
In [SV19] it was shown that the question of whether a given order in a totally definite quaternion algebra has SFC or LFC can be tackled algorithmically.
Theorem 3.5 (Smertnig–Voight).
There exists an algorithm that given an $R$-order $\Lambda$ in a totally definite quaternion algebra over $K$, determines whether $\Lambda$ has SFC (resp. LFC).
This result was used to classify all orders inside totally definite quaternion algebras that have SFC, and of these, all that have LFC (see [SV19, Theorem 1.3]). This classification restricted to SFC and the case of maximal or hereditary orders had previously been given by Vignéras [Vig76], Hallouin–Maire [HM06] and Smertnig [Sme15].
Corollary 3.6.
There exists an algorithm that given a maximal $R$-order $\Lambda$ in a finite-dimensional semisimple $K$-algebra, determines whether $\Lambda$ has SFC (resp. LFC).
Proof.
Let $\mathcal{M}$ be such a maximal $R$-order. Then we can write $\mathcal{M}=\mathcal{M}_{1}\oplus\cdots\oplus\mathcal{M}_{r}$, where each $\mathcal{M}_{i}$ is a maximal $R$-order contained in a simple $K$-algebra $A_{i}$, and each $A_{i}$ either satisfies the Eichler condition or is a totally definite quaternion algebra. Therefore the desired result now follows from Corollary 3.3, Theorem 3.4 and Theorem 3.5. ∎
4. Review of cancellation properties for integral group rings
We recall results concerning SFC and LFC for group rings $\mathbb{Z}[G]$ of a finite group $G$.
Lemma 4.1.
Let $G$ be a finite group and let $H$ be a quotient of $G$. If $\mathbb{Z}[G]$ has SFC (resp. LFC) then $\mathbb{Z}[H]$ has SFC (resp. LFC).
Proof.
The canonical map $\mathbb{Z}[G]\rightarrow\mathbb{Z}[H]$ of $\mathbb{Z}$-orders induces a surjection $\mathbb{Q}[G]\rightarrow\mathbb{Q}[H]$ of $\mathbb{Q}$-algebras, and so the desired result is a consequence of Theorem 3.1. ∎
The finite non-cyclic subgroups of $\mathbb{H}^{\times}$ are called the binary polyhedral groups. These are the generalised quaternion groups $Q_{4n}$ for $n\geq 2$ or one of $\tilde{T}$, $\tilde{O}$ or $\tilde{I}$, the binary tetrahedral, octahedral and icosahedral groups (see [CR87, §51A] and the references therein). It is well-known that a finite group $G$ has a binary polyhedral quotient precisely when the group algebra $\mathbb{Q}[G]$ fails the Eichler condition (see [Nic21a, Proposition 1.5] or [CR87, (51.3)], for example).
We now recall two important results of Swan [Swa83, Theorems I and II].
Theorem 4.2 (Swan).
Let $G$ be a binary polyhedral group. Then $\mathbb{Z}[G]$ has SFC if and only if $\mathbb{Z}[G]$ has LFC if and only if $G$ is one of the following $7$ groups:
$Q_{8},\,Q_{12},\,Q_{16},\,Q_{20},\,\tilde{T},\,\tilde{O},\,\tilde{I}.$
Theorem 4.3 (Swan).
Let $G$ be a binary polyhedral group and let $\mathcal{M}$ be a maximal $\mathbb{Z}$-order in $\mathbb{Q}[G]$. Then $\mathcal{M}$ has LFC if and only if $G$ is one of the following $11$ groups:
$Q_{8},\,Q_{12},\,Q_{16},\,Q_{20},\,Q_{24},\,Q_{28},\,Q_{36},\,Q_{60},\,\tilde{T},\,\tilde{O},\,\tilde{I}.$ (4.1)
Corollary 4.4.
Let $G$ be a finite group and let $\mathcal{M}$ be a maximal $\mathbb{Z}$-order in $\mathbb{Q}[G]$. Then $\mathcal{M}$ has LFC if and only if $G$ has no binary polyhedral quotient other than those in (4.1).
Proof.
Suppose that $G$ has a binary polyhedral quotient $H$ different to those listed in (4.1). Then the image of $\mathcal{M}$ under the canonical projection map $\mathbb{Q}[G]\rightarrow\mathbb{Q}[H]$ is a maximal $\mathbb{Z}$-order, and thus does not have LFC by Theorem 4.3. Hence $\mathcal{M}$ does not have LFC by Theorem 3.1.
Suppose that $G$ has no binary polyhedral quotient other than those listed in (4.1). We can write $\mathcal{M}=\mathcal{M}_{1}\oplus\cdots\oplus\mathcal{M}_{r}$, where each $\mathcal{M}_{i}$ is a maximal $R$-order contained in a simple $K$-algebra $A_{i}$, and each $A_{i}$ either satisfies the Eichler condition or is a totally definite quaternion algebra. By Corollary 3.3, it suffices to show that each $\mathcal{M}_{i}$ has LFC. Fix $i$ with $1\leq i\leq r$. If $A_{i}$ satisfies the Eichler condition then $\mathcal{M}_{i}$ has LFC by Theorem  3.4. So suppose that $A_{i}$ is a totally definite quaternion algebra. Let $\psi_{i}|_{G}:G\rightarrow\mathbb{H}^{\times}$ denote group homomorphism obtained by restricting the composition $\psi_{i}:\mathbb{Q}[G]\twoheadrightarrow A_{i}\hookrightarrow\mathbb{H}$. Let $H_{i}$ and $N_{i}$ denote the image and kernel of $\psi_{i}|_{G}$, respectively. Then the elements of $H_{i}$ span $\mathbb{H}$ over $\mathbb{R}$ and so $H_{i}$ must be non-abelian since $\mathbb{H}$ is a noncommutative ring. Thus $H_{i}$ is a binary polyhedral quotient of $G$ and so by hypothesis $H_{i}$ is isomorphic to one of the groups listed in (4.1). Let $e_{N_{i}}=|N_{i}|^{-1}\sum_{n\in N_{i}}n$ denote the central idempotent of $\mathbb{Q}[G]$ attached to $N_{i}$. Then
$\mathbb{Q}[G]=\mathbb{Q}[G]e_{N_{i}}\oplus\mathbb{Q}[G](1-e_{N_{i}})$
and since $\psi_{i}|_{\mathbb{Q}[G](1-e_{N_{i}})}=0$ it follows that $A_{i}$ is a direct summand of $\mathbb{Q}[G]e_{N_{i}}\cong\mathbb{Q}[H_{i}]$. Therefore $\mathcal{M}_{i}$ is a direct summand of a maximal $\mathbb{Z}$-order $\mathcal{M}_{i}^{\prime}$ containing $\mathbb{Z}[H_{i}]$. Then $\mathcal{M}_{i}^{\prime}$ has LFC by Theorem 4.3 and thus $\mathcal{M}_{i}$ also has LFC by Corollary 3.3. ∎
The following two results are [Swa83, Theorems 13.7 and 15.5]. Though the latter result is stated in loc. cit. for LFC only, it is immediate from the proof that it also holds for SFC; the authors are grateful to John Nicholson for bringing this to their attention.
Theorem 4.5 (Swan).
The group ring $\mathbb{Z}[\tilde{T}^{m}\times\tilde{I}^{n}]$ has LFC for all $m,n\geq 0$.
Theorem 4.6 (Swan).
Let $G$ and $H$ be finite groups such that $\mathbb{Z}[G]$ has SFC (resp. LFC), $\mathbb{Q}[H]$ satisfies the Eichler condition, and $H$ has no subgroup of index $2$. Then $\mathbb{Z}[G\times H]$ has SFC (resp. LFC).
The following results on the failure of SFC are [Swa83, Theorem 15.1] and a summary of the main results of Chen [Che86]. Though the latter result is stated for LFC in loc. cit., a close examination of the proofs shows that it also holds for SFC.
Theorem 4.7 (Swan).
The group ring $\mathbb{Z}[Q_{8}\times C_{2}]$ fails SFC.
Theorem 4.8 (Chen).
The group ring $\mathbb{Z}[G]$ fails SFC if $G$ is one of the following $5$ groups:
$Q_{12}\times C_{2},\,Q_{16}\times C_{2},\,Q_{20}\times C_{2},\,G_{(36,7)},G_{(100,7)}.$
[…]
Theorem 4.9 (Nicholson).
Let $G$ be a finite group with quotient $H$ such that $m_{\mathbb{H}}(G)=m_{\mathbb{H}}(H)$ and suppose that $K_{1}(\mathbb{Z}[H])$ is represented by units. Then $\mathbb{Z}[G]$ has SFC if and only if $\mathbb{Z}[H]$ has SFC.
Theorem 4.10 (Nicholson).
Let $G$ be a finite group with binary polyhedral quotient $H$ such that $m_{\mathbb{H}}(G)=m_{\mathbb{H}}(H)$. Then $\mathbb{Z}[G]$ has SFC if and only if $\mathbb{Z}[H]$ has SFC.
Theorem 4.11 (Nicholson).
If $G$ is a finite group with $m_{\mathbb{H}}(G)\leq 1$ then $\mathbb{Z}[G]$ has SFC.
For a finite group $G$ and a field $\mathbb{F}$, let $r_{\mathbb{F}}(G)$ denote the number of irreducible $\mathbb{F}$-representations of $G$, and let $SK_{1}(\mathbb{Z}[G])=\ker(K_{1}(\mathbb{Z}[G])\rightarrow K_{1}(\mathbb{Q}[G]))$; this is a finite abelian group by a theorem of Bass [CR87, (45.20)]. We say that $K_{1}(\mathbb{Z}[G])$ is represented by trivial units if the canonical map ±$\pm G\rightarrow K_{1}(\mathbb{Z}[G])$ is surjective.
As explained in [Nic21a, p. 324], the following result can be deduced from results of Bass [Bas65, Corollary 6.3] (see also [CR87, (45.21)(ii)]) and Wall [Wal74, Proposition 6.5] (see also [CR87, (46.4)]), and in certain situations, can be used to apply Theorem 4.9.
Proposition 4.12.
Let $G$ be a finite group such that $SK_{1}(\mathbb{Z}[G])=0$ and $r_{\mathbb{R}}(G)=r_{\mathbb{Q}}(G)$. Then $K_{1}(\mathbb{Z}[G])$ is represented by trivial units.
5. Fibre products and reduction criteria for SFC
Let $R$ be a Dedekind domain whose quotient field $K$ is a number field. Let $A$ be a finite-dimensional semisimple $K$-algebra and let $\Lambda$ be an $R$-order in $A$.
5.1. Fibre products
Suppose that there is a fibre product
${\Lambda}$${\Lambda_{1}}$${\Lambda_{2}}$${\overline{\Lambda}.}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{g_{1}}$$\scriptstyle{g_{2}}$ (5.1)
in which $\Lambda_{1}$ and $\Lambda_{2}$ are $R$-orders in finite-dimensional semisimple $K$-algebras $A_{1}$, and $A_{2}$, respectively, and where $\overline{\Lambda}$ is an $R$-torsion $R$-algebra. In particular, this means that there are $R$-algebra isomorphisms
$\Lambda\cong\{(x_{1},x_{2})\in\Lambda_{1}\oplus\Lambda_{2}\mid g_{1}(x_{1})=g_{2}(x_{2})\}\quad\textrm{ and }\quad A\cong A_{1}\oplus A_{2},$ (5.2)
which we take to be identifications.
Example 5.1.
Let $I$ and $J$ be two-sided ideals of $\Lambda$ such that $I\cap J=0$ and $K(I+J)=A$. Let $\Lambda_{1}=\Lambda/I$, let $\Lambda_{2}=\Lambda/J$ and let $\overline{\Lambda}=\Lambda/(I+J)$. Then $\overline{\Lambda}$ is $R$-torsion and there is a fibre product as in (5.1) such that each of $\pi_{1},\pi_{2},g_{1},g_{2}$ is the relevant canonical surjection.
5.2. Review of results of Reiner–Ullom
We now recall [RU74, Lemma 4.20], which is crucial to all of the results in this section.
Proposition 5.2 (Reiner–Ullom).
Given a fibre product as in (5.1), assume that at least one of $g_{1}$ and $g_{2}$ is surjective. For $u\in\overline{\Lambda}^{\times}$, define
$M(u):=\{(x_{1},x_{2})\in\Lambda_{1}\oplus\Lambda_{2}\mid g_{1}(x_{1})u=g_{2}(x_{2})\}.$
Then the following statements hold.
  1. (i)
    Each $M(u)$ is a locally free $\Lambda$-lattice of rank $1$ such that
    $\Lambda_{i}\otimes_{\Lambda}M(u)\cong\Lambda_{i}\cdot M(u)=\Lambda_{i}\textrm{for }i=1,2.$
  2. (ii)
    For $u,u^{\prime}\in\overline{\Lambda}^{\times}$, we have
    $M(u)\oplus M(u^{\prime})\cong\Lambda\oplus M(uu^{\prime})\cong\Lambda\oplus M(u^{\prime}u).$
  3. (iii)
    Let $\Lambda^{\prime}$ be a maximal $R$-order in $A$ containing $\Lambda$. Then
    $\Lambda^{\prime}\otimes_{\Lambda}M(u)\cong\Lambda^{\prime}\cdot M(u)=\Lambda^{\prime}.$
  4. (iv)
    Let $X$ be any locally free $\Lambda$-lattice such that $\Lambda_{i}\otimes_{\Lambda}X\cong\Lambda_{i}$ for $i=1,2$. Then $X\cong M(u)$ for some $u\in\overline{\Lambda}^{\times}$.
  5. (v)
    For $u,u^{\prime}\in\overline{\Lambda}^{\times}$, we have $M(u)\cong M(u^{\prime})$ if and only if $u^{\prime}\in g_{1}(\Lambda_{1}^{\times})\cdot u\cdot g_{2}(\Lambda_{2}^{\times})$.
Corollary 5.3.
In the setting of Proposition 5.2, the following statements hold.
  1. (i)
    The map $\partial:\overline{\Lambda}^{\times}\rightarrow\mathrm{Cl}(\Lambda)$ defined by $\partial(u)=[M(u)]_{\mathrm{st}}$ is a homomorphism.
  2. (ii)
    For $u\in\overline{\Lambda}^{\times}$, we have that $M(u)\cong\Lambda$ if and only if $u\in W$, where
    $W:=\{w_{1}w_{2}\mid w_{i}\in g_{i}(\Lambda_{i}^{\times}),i=1,2\}.$
  3. (iii)
    For $u\in\overline{\Lambda}^{\times}$, we have that $M(u)$ is stably free over $\Lambda$ if and only if $u\in\ker(\partial)$.
  4. (iv)
    We have a containment of sets $W\subseteq\ker(\partial)$.
Proof.
Parts (i) and (ii) follow from Proposition 5.2 parts (ii) and (v), respectively. Part (iii) follows from the definition of locally free class group given in §2.5. Part (iv) follows from parts (ii) and (iii). ∎
5.3. A reduction criterion for SFC
The following underpins many of our results.
Proposition 5.4.
Assume the setting of Corollary 5.3. Then $\Lambda$ has SFC if and only if both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC and $W=\ker(\partial)$.
Proof.
Suppose that $\Lambda$ has SFC. The maps $K\otimes_{R}\pi_{i}:A\rightarrow A_{i}$ are surjective for $i=1,2$ by (5.2). Thus both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC by Theorem 3.1. Let $u\in\ker(\partial)$. Then $\partial(u)=[M(u)]_{\mathrm{st}}=0$ in $\mathrm{Cl}(\Lambda)$ and so $M(u)\cong\Lambda$ since $\Lambda$ has SFC. Thus $u\in W$ by Corollary 5.3 (ii). Hence $\ker(\partial)\subseteq W$. But the reverse containment also holds by Corollary 5.3 (iv) and so $W=\ker(\partial)$.
Suppose conversely that both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC and that $W=\ker(\partial)$. Recall from §3.1 that in order to show that $\Lambda$ has SFC, it suffices to show that every stably free $\Lambda$-lattice of rank $1$ is in fact free of rank $1$. So suppose that $X$ is a stably free $\Lambda$-lattice of rank $1$, that is, $X\in g(\Lambda)$ and $[X]_{\mathrm{st}}=0$ in $\mathrm{Cl}(\Lambda)$. Then for $i=1,2$ we have $[\Lambda_{i}\otimes_{\Lambda}X]_{\mathrm{st}}=0$ in $\mathrm{Cl}(\Lambda_{i})$ and so $\Lambda_{i}\otimes_{\Lambda}X\cong\Lambda_{i}$ since $\Lambda_{i}$ has SFC. Hence $X\cong M(u)$ for some $u\in\overline{\Lambda}^{\times}$ by Proposition 5.2 (iv). Since $X$ is stably free, Corollary 5.3 (iii) implies that $u\in\ker(\partial)$. Thus since $W=\ker(\partial)$, Corollary 5.3 (ii) implies that $M(u)\cong\Lambda$. Therefore $X\cong M(u)\cong\Lambda$. ∎
Corollary 5.5.
Assume the setting of Corollary 5.3. Suppose that both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC and that $W$ is a group. Let $N$ be a normal subgroup of $\ker(\partial)$ contained in $W$ and let $S$ be a generating subset of $\ker(\partial)/N$. Then the following assertions are equivalent.
  1. (i)
    $\Lambda$ has SFC.
  2. (ii)
    For every $\overline{u}\in\ker(\partial)/N$ and every lift $u\in\ker(\partial)$ we have $M(u)\cong\Lambda$.
  3. (iii)
    For every $\overline{u}\in S$ there exists a lift $u\in\ker(\partial)$ such that $M(u)\cong\Lambda$.
Proof.
We shall show that (i) $\implies$ (ii) $\implies$ (iii) $\implies$ (i). Suppose (i) holds. Then $W=\ker(\partial)$ by Proposition 5.4. Thus $M(u)\cong\Lambda$ for every $u\in\ker(\partial)$ by Corollary 5.3 (ii) and so, in particular, assertion (ii) holds. Clearly (ii) implies (iii). Suppose (iii) holds. By Proposition 5.4, to show that (i) holds, it suffices to show that $\ker(\partial)=W$. Let $v\in\ker(\partial)$ and let $\overline{v}$ denote its image in $\ker(\partial)/N$. Then there exist (not necessarily distinct) elements $\overline{s}_{1},\ldots,\overline{s}_{r}\in S$ such that $\overline{v}=\overline{s}_{1}\cdots\overline{s}_{r}$. By hypothesis, there exist lifts $s_{1},\ldots,s_{r}\in\ker(\partial)$ such that $M(s_{1})\cong\cdots\cong M(s_{r})\cong\Lambda$, and so $s_{1},\ldots,s_{r}\in W$ by Corollary 5.3 (ii). Moreover, there exists $n\in N$ such that $v=s_{1}\cdots s_{r}n$. Since $W$ is a group and $N\leq W$, we conclude that $v\in W$. Therefore $\ker(\partial)\subseteq W$, and hence $\ker(\partial)=W$ by Corollary 5.3 (iv). ∎
5.4. A criterion in terms of sizes of locally free class groups
The following is a variant of [CR87, (49.30)], which itself is based on [RU74, §4, §5].
Proposition 5.6.
Assume the setting of Corollary 5.3. Then there is an exact sequence
$1\longrightarrow\ker(\partial)\longrightarrow\overline{\Lambda}^{\times}\stackrel{{\scriptstyle\partial}}{{\longrightarrow}}\mathrm{Cl}(\Lambda)\longrightarrow\mathrm{Cl}(\Lambda_{1})\oplus\mathrm{Cl}(\Lambda_{2})\longrightarrow 0$ (5.3)
Proof.
Since $\overline{\Lambda}$ is semilocal, the canonical map $\pi:\overline{\Lambda}^{\times}\rightarrow K_{1}(\overline{\Lambda})$ is surjective by [CR87, (40.31)]. Moreover, by [CR87, (49.27)] there is an exact sequence
$K_{1}(\Lambda_{1})\times K_{1}(\Lambda_{2})\longrightarrow K_{1}(\overline{\Lambda})\stackrel{{\scriptstyle\partial^{\prime}}}{{\longrightarrow}}\mathrm{Cl}(\Lambda)\longrightarrow\mathrm{Cl}(\Lambda_{1})\oplus\mathrm{Cl}(\Lambda_{2})\longrightarrow 0.$ (5.4)
As shown in the proof of loc. cit. $\partial^{\prime}$ can be defined by $\partial=\partial^{\prime}\circ\pi$ and taking lifts via $\pi$. (Note that the map $\partial$ here is different to that of loc. cit.) Thus $\mathrm{image}(\partial)=\mathrm{image}(\partial^{\prime})$. Therefore the desired result follows from the existence and exactness of (5.4). ∎
Corollary 5.7.
Assume the setting of Corollary 5.3. Then we have
$|W|\leq|\ker(\partial)|=|\mathrm{Cl}(\Lambda_{1})||\mathrm{Cl}(\Lambda_{2})||\overline{\Lambda}^{\times}||\mathrm{Cl}(\Lambda)|^{-1}.$ (5.5)
Moreover, $\Lambda$ has SFC if and only if both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC and (5.5) is an equality.
Proof.
Proposition 5.6 shows that $|\ker(\partial)|=|\mathrm{Cl}(\Lambda_{1})||\mathrm{Cl}(\Lambda_{2})||\overline{\Lambda}^{\times}||\mathrm{Cl}(\Lambda)|^{-1}$. The result now follows from Corollary 5.3 (iv) and Proposition 5.4. ∎
Remark 5.8.
The locally free class groups in Proposition 5.6 and Corollary 5.7 can be replaced by the corresponding kernel groups as defined in [CR87, (49.33)]; this follows from Proposition 5.2 (iii) and [CR87, (49.39)]. Thus (5.3) becomes
$1\longrightarrow\ker(\partial)\longrightarrow\overline{\Lambda}^{\times}\stackrel{{\scriptstyle\partial}}{{\longrightarrow}}D(\Lambda)\longrightarrow D(\Lambda_{1})\oplus D(\Lambda_{2})\longrightarrow 0,$
and so (5.5) becomes $|W|\leq|\ker(\partial)|=|D(\Lambda_{1})||D(\Lambda_{2})||\overline{\Lambda}^{\times}||D(\Lambda)|^{-1}$. Both locally free class groups and kernel groups can be calculated by using the algorithm of [BB06, §3].
Remark 5.9.
For any associative ring $\mathfrak{R}$ with identity, we have a short exact sequence
$1\longrightarrow 1+J(\mathfrak{R})\longrightarrow\mathfrak{R}^{\times}\longrightarrow\left(\mathfrak{R}/J(\mathfrak{R})\right)^{\times}\longrightarrow 1,$
where $J(\mathfrak{R})$ denotes the Jacobson radical of $\mathfrak{R}$ (see [CR81, Exercise 5.2], for example). In particular, if $\mathfrak{R}$ is finite then |×$|\mathfrak{R}^{\times}|=|J(\mathfrak{R})||(\mathfrak{R}/J(\mathfrak{R}))^{\times}|$. If $\mathfrak{R}$ is a finite-dimensional algebra over a finite field, then $J(\mathfrak{R})$ can be computed using the algorithm of [FR85].
5.5. A criterion assuming the Eichler condition for one component
The following is a specialisation of [Swa80, Theorem 10.2].
Theorem 5.10 (Swan).
Let $\mathfrak{I}$ be a two-sided ideal of $\Lambda$ such that $\Lambda/\mathfrak{I}$ is $R$-torsion and let $f:\Lambda\rightarrow\Lambda/\mathfrak{I}$ be the canonical quotient map. Assume that $A$ satisfies the Eichler condition. Then $f(\Lambda^{\times})$ contains the commutator subgroup $[(\Lambda/\mathfrak{I})^{\times},(\Lambda/\mathfrak{I})^{\times}]$.
This leads to the following result whose corollary is useful for computational purposes.
Proposition 5.11.
Let $I$ and $J$ be two-sided ideals of $\Lambda$ such that $I\cap J=0$ and $K(I+J)=A$. Let $\Lambda_{1}=\Lambda/I$, let $\Lambda_{2}=\Lambda/J$ and let $\overline{\Lambda}=\Lambda/(I+J)$. Assume that $A_{1}:=K\Lambda_{1}\cong KJ$ satisfies the Eichler condition. For $i=1,2$, let $g_{i}:\Lambda_{i}\rightarrow\overline{\Lambda}$ be the canonical surjection. Let $\partial$ and $W$ be as in Corollary 5.3. Then $\overline{\Lambda}$ is finite and we have
$[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]\trianglelefteq g_{1}(\Lambda_{1}^{\times})\trianglelefteq W\trianglelefteq\ker(\partial)\trianglelefteq\overline{\Lambda}^{\times},$
where each term is a normal subgroup of $\overline{\Lambda}^{\times}$ with finite abelian quotient.
Proof.
The hypotheses imply that we are in the setting of Example 5.1. In particular, $\overline{\Lambda}$ is $R$-torsion and hence finite. Moreover, every subgroup of $\overline{\Lambda}^{\times}$ containing $[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]$ is necessarily normal in $\overline{\Lambda}^{\times}$ and has finite abelian quotient. Theorem 5.10 applied with $f=g_{1}$ implies that $[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]\trianglelefteq g_{1}(\Lambda_{1}^{\times})$. Since $W=g_{1}(\Lambda_{1}^{\times})g_{2}(\Lambda_{2}^{\times})$ by definition and $g_{1}(\Lambda_{1}^{\times})$ is normal in $\overline{\Lambda}^{\times}$, we have that $W$ is in fact a subgroup of $\overline{\Lambda}^{\times}$ (a priori, it is only a subset), and that $g_{1}(\Lambda_{1}^{\times})\trianglelefteq W$. The remaining containments follow from Corollary 5.3. ∎
Corollary 5.12.
Assume the setting of Proposition 5.11. Then $W$ is a group. Suppose that $N$ is a subgroup of $W$ containing $[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]$. Then $N\trianglelefteq\ker(\partial)$ and $\ker(\partial)/N$ is a finite abelian group. Let $S$ be a generating subset of $\ker(\partial)/N$. Suppose that $\Lambda_{2}$ has SFC. Then the following assertions are equivalent.
  1. (i)
    $\Lambda$ has SFC.
  2. (ii)
    For every $\overline{u}\in\ker(\partial)/N$ and every lift $u\in\ker(\partial)$ we have $M(u)\cong\Lambda$.
  3. (iii)
    For every $\overline{u}\in S$ there exists a lift $u\in\ker(\partial)$ such that $M(u)\cong\Lambda$.
Proof.
The assertions that $W$ is a group, $N\trianglelefteq\ker(\partial)$ and $\ker(\partial)/N$ is a finite abelian group all follow from Proposition 5.11. Since $A_{1}$ satisfies the Eichler condition, $\Lambda_{1}$ has SFC by Theorem 3.4. Moreover, $\Lambda_{2}$ satisfies SFC by hypothesis. Thus assertions (i), (ii) and (iii) are equivalent by Corollary 5.5. ∎
6. Applications to integral group rings
In this section, we consider applications of Corollary 5.7 to integral group rings.
6.1. A fibre product for integral group rings
We first record the following well-known instance of a fibre product. Let $G$ be a finite group, let $N$ be a normal subgroup and let $H=G/N$. Let $n=|N|$ and let $\mathrm{Tr}_{N}=\sum_{n\in N}n\in\mathbb{Z}[G]$ be the trace element associated to $N$. Let $\pi_{1}:\mathbb{Z}[G]\rightarrow\mathbb{Z}[H]$ be the canonical projection and let $I=\ker(\pi_{1})$. Let $J=\mathrm{Tr}_{N}\mathbb{Z}[G]$, let $\Gamma=\mathbb{Z}[G]/J$ and let $\pi_{2}:\mathbb{Z}[G]\rightarrow\mathbb{Z}[G]/J$ be the canonical projection. Then $I\cap J=0$ and $I+J=n\mathbb{Z}[G]$. Thus, by Example 5.1 we have a fibre product
${\mathbb{Z}[G]}$${\mathbb{Z}[H]}$${\Gamma}$${(\mathbb{Z}/n\mathbb{Z})[H],}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{g_{1}}$$\scriptstyle{g_{2}}$ (6.1)
where each of $\pi_{1},\pi_{2},g_{1},g_{2}$ is the relevant canonical surjection. Let
$W_{G,N}:=\{w_{1}w_{2}\mid w_{1}\in g_{1}(\mathbb{Z}[H]^{\times}),w_{2}\in g_{2}(\Gamma^{\times})\}.$
Then Corollary 5.7 immediately implies the following result.
Proposition 6.1.
Assume the above setting and notation. Then
$|W_{G,N}|\leq\frac{|\mathrm{Cl}(\mathbb{Z}[H])||\mathrm{Cl}(\mathbb{Z}[\Gamma])||(\mathbb{Z}/n\mathbb{Z})[H]^{\times}|}{|\mathrm{Cl}(\mathbb{Z}[G])|}.$ (6.2)
Moreover, $\mathbb{Z}[G]$ has SFC if and only if both $\mathbb{Z}[H]$ and $\Gamma$ have SFC and (6.2) is an equality.
Remark 6.2.
The locally free class groups in Proposition 6.1 (and in Propositions 6.4 and 6.10 below) can be replaced by the corresponding kernel groups (see Remark 5.8).
Remark 6.3.
Let $p$ be a prime number and suppose that $n=p$ in Proposition 6.1. Then $(\mathbb{Z}/n\mathbb{Z})[H]=\mathbb{F}_{p}[H]$. Thus left multiplication by an element of $\mathbb{F}_{p}[H]^{\times}$ induces an $\mathbb{F}_{p}$-linear automorphism of $\mathbb{F}_{p}[H]$. Together with the choice of $H$ as an $\mathbb{F}_{p}$-basis of $\mathbb{F}_{p}[H]$, this determines an explicit embedding of $\mathbb{F}_{p}[H]^{\times}$ into $\mathrm{GL}_{h}(\mathbb{F}_{p})$, where $h=|H|$. This allows computations in $\mathbb{F}_{p}[H]^{\times}$ to be performed efficiently on a computer algebra system. By contrast, in the more general setting of Corollary 5.7, it can be difficult to perform computations in $\overline{\Lambda}^{\times}$ (see §11.3).
6.2. Integral group rings of the form $\mathbb{Z}[H\times C_{2}]$
We will need the following result.
Proposition 6.4.
Let $H$ be a finite group such that $\mathbb{Z}[H]$ has SFC. Let $V_{H}$ be the image of the canonical projection $\mathbb{Z}[H]^{\times}\rightarrow\mathbb{F}_{2}[H]^{\times}$. Then
$|V_{H}|\leq\frac{|\mathrm{Cl}(\mathbb{Z}[H])|^{2}|\mathbb{F}_{2}[H]^{\times}|}{|\mathrm{Cl}(\mathbb{Z}[H\times C_{2}])|},$
and this inequality is an equality if and only if $\mathbb{Z}[H\times C_{2}]$ has SFC.
Proof.
Let $G=H\times C_{2}$ and let $N=\{1\}\times C_{2}\leq G$. Then in the notation of §6.1, we have $n=2$, $\Gamma\simeq\mathbb{Z}[H]$ and $V_{H}=W_{G,N}$. Thus the desired result now follows immediately from Proposition 6.1. ∎
Example 6.5.
Let $Q_{8}$ denote the quaternion group of order $8$. Swan [Swa83, Theorem 15.1] showed that $\mathbb{Z}[Q_{8}\times C_{2}]$ does not have SFC. We now give an alternative proof of this result using Proposition 6.4 and the following facts:
  1. (i)
    $\mathbb{Z}[Q_{8}]$ has SFC ([Swa83, Theorem I]),
  2. (ii)
    $\mathbb{Z}[Q_{8}]^{\times}=\pm Q_{8}$ ([Swa83, Lemma 15.3]),
  3. (iii)
    $|\mathrm{Cl}(\mathbb{Z}[Q_{8}])|=2$ ([Swa83, Theorem III]),
  4. (iv)
    $|\mathrm{Cl}(\mathbb{Z}[Q_{8}\times C_{2}])|=2^{4}$ ([Swa83, Theorem 16.5]), and
  5. (v)
    $|\mathbb{F}_{2}[Q_{8}]^{\times}|=2^{7}$ (since $\mathbb{F}_{2}[Q_{8}]$ is a local ring).
Fact (ii) implies that $V_{Q_{8}}=Q_{8}\leq\mathbb{F}_{2}[Q_{8}]^{\times}$. Hence
$2^{3}=|Q_{8}|=|V_{Q_{8}}|<|\mathrm{Cl}(\mathbb{Z}[Q_{8}])|^{2}|\mathrm{Cl}(\mathbb{Z}[Q_{8}\times C_{2}])|^{-1}|\mathbb{F}_{2}[Q_{8}]^{\times}|=2^{2}\cdot 2^{-4}\cdot 2^{7}=2^{5},$
and so $\mathbb{Z}[Q_{8}\times C_{2}]$ does not have SFC. Note that the quantities in (iii) and (iv) can instead both be calculated on a computer by using the algorithm of [BB06, §3].
Let $\tilde{T}$ denote the binary tetrahedral group. Then $\tilde{T}\cong\mathrm{SL}_{2}(\mathbb{F}_{3})\cong G_{(24,3)}$.
Theorem 6.6.
The integral group ring $\mathbb{Z}[\tilde{T}\times C_{2}]$ has SFC.
Proof.
The algorithm of [BB06, §3] shows that $|\mathrm{Cl}(\mathbb{Z}[\tilde{T}])|=2$ and $|\mathrm{Cl}(\mathbb{Z}[\tilde{T}\times C_{2}])|=2^{6}$. By Remark 5.9, we have $|\mathbb{F}_{2}[\tilde{T}]^{\times}|=|\mathbb{F}_{2}[\tilde{T}]||(\mathbb{F}_{2}[\tilde{T}]/J(\mathbb{F}_{2}[\tilde{T}]))^{\times}|$, where $J(\mathbb{F}_{2}[\tilde{T}])$ denotes the Jacobson radical of $\mathbb{F}_{2}[\tilde{T}]$. By [MS22, Theorem 2.5] we have $\mathbb{F}_{2}[\tilde{T}]/J(\mathbb{F}_{2}[\tilde{T}])\cong\mathbb{F}_{2}\oplus\mathbb{F}_{4}$ (this can also be verified using the algorithm of [FR85]). Hence $|J(\mathbb{F}_{2}[\tilde{T}])|=2^{24-3}=2^{21}$. Moreover, $(\mathbb{F}_{2}[\tilde{T}]/J(\mathbb{F}_{2}[\tilde{T}]))^{\times}\cong(\mathbb{F}_{2}\oplus\mathbb{F}_{4})^{\times}\cong C_{3}$. Therefore $|\mathbb{F}_{2}[\tilde{T}]^{\times}|=3\cdot 2^{21}$.
The integral group ring $\mathbb{Z}[\tilde{T}]$ has SFC by [Swa83, Theorem I]. Let $V=V_{\tilde{T}}$ be the image of the canonical projection $\pi:\mathbb{Z}[\tilde{T}]^{\times}\rightarrow\mathbb{F}_{2}[\tilde{T}]^{\times}$. Then by Proposition 6.4 and the above calculations we have
$|V|\leq\frac{|\mathrm{Cl}(\mathbb{Z}[\tilde{T}])|^{2}|\mathbb{F}_{2}[\tilde{T}]^{\times}|}{|\mathrm{Cl}(\mathbb{Z}[\tilde{T}\times C_{2}])|}=3\cdot 2^{17},$
with equality if and only if $\mathbb{Z}[\tilde{T}\times C_{2}]$ has SFC.
To prove equality, we consider the following units of integral groups rings as introduced by Ritter and Segal in [RS91]. For $g,h\in\tilde{T}$, let $d=\operatorname{ord}(g)$ and define
$u_{g,h}$ $=1+(1-g)h(1+g+\dotsb+g^{d-1}),$
$u^{\prime}_{g,h}$ $=1+(1+g+\dotsb+g^{d-1})h(1-g).$
Then $u_{g,h}$ and $u^{\prime}_{g,h}$ are elements of $\mathbb{Z}[\tilde{T}]^{\times}$. Let
$U=\langle g,\pi(u_{g,h}),\pi(u^{\prime}_{g,h})\mid g,h\in\tilde{T}\rangle$
and note that this is a subgroup of $V$. By Remark 6.3, we obtain an explicit embedding of $\mathbb{F}_{2}[\tilde{T}]^{\times}$ into $\mathrm{GL}_{24}(\mathbb{F}_{2})$. Since a calculation on a computer algebra system such as Magma [BCP97] or Oscar [OSC24, DEF+24] shows that the image of $U$ under this embedding is a subgroup of $\mathrm{GL}_{24}(\mathbb{F}_{2})$ of order $3\cdot 2^{17}$, we must have $U=V$ and $|V|=3\cdot 2^{17}$. Therefore we conclude that $\mathbb{Z}[\tilde{T}\times C_{2}]$ has SFC. ∎
Remark 6.7.
Using Theorem 6.6 as an input, Nicholson [Nic24, Theorem 1.7] has shown that the integral group ring $\mathbb{Z}[\tilde{T}\times C_{2}]$ in fact has LFC.
Lemma 6.8.
We have $SK_{1}(\mathbb{Z}[\tilde{T}\times C_{2}])=0$.
Proof.
By [Oli81, Theorem 3], for any finite group $G$ and any prime $p$, the Sylow $p$-subgroup of $SK_{1}(\mathbb{Z}[G])$ is generated by induction from the $p$-elementary subgroups of $G$. If $H$ is any abelian subgroup of $\tilde{T}\times C_{2}$ then $H\cong C_{2}\times C_{2}$, $C_{2}\times C_{4}$, $C_{2}\times C_{6}$ or $H$ is cyclic and so $SK_{1}(\mathbb{Z}[H])=0$ by [ADOS87, Theorem 4.9]. If $H$ is any non-abelian elementary subgroup of $\tilde{T}\times C_{2}$ then $H\cong Q_{8}$ or $Q_{8}\times C_{2}$ and so $SK_{1}(\mathbb{Z}[H])=0$ by [Oli81, Theorem 4] and [Oli88, Example 9.9(ii)], respectively. Therefore $SK_{1}(\mathbb{Z}[\tilde{T}\times C_{2}])=0$. ∎
We now have the following consequence of Theorem 6.6, which has also been shown independently by Nicholson [Nic24, Proposition 4.14] .
Corollary 6.9.
Let $G$ be a finite group with quotient isomorphic to $\tilde{T}\times C_{2}$ and with $m_{\mathbb{H}}(G)=2$. Then $\mathbb{Z}[G]$ has SFC.
Proof.
Let $H=\tilde{T}\times C_{2}$. Then $SK_{1}(\mathbb{Z}[H])=0$ by Lemma 6.8 and it is straightforward to deduce from the character table of $H$ that $r_{\mathbb{R}}(H)=r_{\mathbb{Q}}(H)=10$. Thus $K_{1}(\mathbb{Z}[H])$ is represented by (trivial) units by Proposition 4.12. We also have that $m_{\mathbb{H}}(H)=2$, which can be deduced from the character table of $H$ along with its Frobenius-Schur indicators. Thus the desired result now follows from Theorems 4.9 and 6.6. ∎
6.3. A useful reduction criterion for certain integral group rings
The following reduction criterion only uses the sizes of certain locally free class groups, and thus can be applied relatively easily.
Proposition 6.10.
Let $G$ be a finite group with a normal subgroup $N$ of size $2$ and let $H=G/N$. Let $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$. Suppose that $\mathbb{Z}[H\times C_{2}]$ has SFC and that
$\frac{|\mathrm{Cl}(\mathbb{Z}[H\times C_{2}])|}{|\mathrm{Cl}(\mathbb{Z}[H])|}=\frac{|\mathrm{Cl}(\mathbb{Z}[G])|}{|\mathrm{Cl}(\Gamma)|}.$ (6.3)
Then $\mathbb{Z}[G]$ has SFC if and only if $\Gamma$ has SFC.
Proof.
Since $\mathbb{Z}[H\times C_{2}]$ has SFC by hypothesis, so does $\mathbb{Z}[H]$ by Lemma 4.1. By Proposition 6.4 and (6.3) we have
$|V_{H}|=\frac{|\mathrm{Cl}(\mathbb{Z}[H])|^{2}|\mathbb{F}_{2}[H]^{\times}|}{|\mathrm{Cl}(\mathbb{Z}[H\times C_{2}])|}=\frac{|\mathrm{Cl}(\mathbb{Z}[H])||\mathbb{F}_{2}[H]^{\times}||\mathrm{Cl}(\Gamma)|}{|\mathrm{Cl}(\mathbb{Z}[G])|}.$ (6.4)
By Proposition 6.1 and (6.4) we have
$|W_{G,N}|\leq\frac{|\mathrm{Cl}(\mathbb{Z}[H])||\mathrm{Cl}(\Gamma)||\mathbb{F}_{2}[H]^{\times}|}{|\mathrm{Cl}(\mathbb{Z}[G])|}=|V_{H}|,$
and that $\mathbb{Z}[G]$ has SFC if and only if $\Gamma$ has SFC and this inequality is in fact an equality. But $V_{H}\subseteq W_{G,N}$ by the definitions. Therefore $|W_{G,N}|=|V_{H}|$ and so $\mathbb{Z}[G]$ has SFC if and only if $\Gamma$ has SFC. ∎
7. Criteria for freeness and stable freeness
7.1. Setup
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$. By [Rei03, (10.4)] there exists a (not necessarily unique) maximal $\mathcal{O}$-order $\mathcal{M}$ in $A$ containing $\Lambda$. Let $\mathfrak{f}$ be any proper full two-sided ideal of $\mathcal{M}$ that is contained in $\Lambda$.
7.2. A criterion for freeness
We recall some results from [BHJ22, §4]. Define
$\pi:(\mathcal{M}/\mathfrak{f})^{\times}\longrightarrow{(\mathcal{M}/\mathfrak{f})^{\times}}/{(\Lambda/\mathfrak{f})^{\times}}$ (7.1)
to be the map induced by the canonical projection, where the codomain is the collection of left cosets of $(\Lambda/\mathfrak{f})^{\times}$ in $(\mathcal{M}/\mathfrak{f})^{\times}$. Note that $(\Lambda/\mathfrak{f})^{\times}$ is a subgroup of $(\mathcal{M}/\mathfrak{f})^{\times}$, but is not necessarily a normal subgroup, and so $\pi$ is only a map of sets in general.
For $\eta\in\mathcal{M}$ we write $\overline{\eta}$ for its image in $\mathcal{M}/\mathfrak{f}$.
Proposition 7.1 ([BHJ22, Proposition 4.3]).
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Let $u\in\mathcal{M}^{\times}$ and let $\alpha=u\beta$. Then $\overline{\alpha},\overline{\beta},\overline{u}\in(\mathcal{M}/\mathfrak{f})^{\times}$ and the following are equivalent:
  1. (i)
    $X=\Lambda\alpha$,
  2. (ii)
    $\Lambda\alpha+\mathfrak{f}=\Lambda$,
  3. (iii)
    $\overline{\alpha}\in(\Lambda/\mathfrak{f})^{\times}$,
  4. (iv)
    $\pi(\overline{\beta})=\pi(\overline{u^{-1}})$,
  5. (v)
    $\alpha\in X$ and $X$ is locally free over $\Lambda$.
Remark 7.2.
[BHJ22, Proposition 4.3 and Theorem 4.5] only assumed that $\beta\in\mathcal{M}$ but should have assumed that $\beta\in\mathcal{M}\cap A^{\times}$, or equivalently, that $\beta\in\mathcal{M}$ and that $X$ is a full left ideal of $\Lambda$, that is, $KX=A$. Note that any locally free left ideal of $\Lambda$ is necessarily full. Moreover, if $\Lambda\cap V\neq\mathfrak{f}\cap V$ for all non-trivial $K$-vector subspaces $V$ of $A$, then the condition $X+\mathfrak{f}=\Lambda$ also implies that $X$ is full. The strengthened assumption is needed to ensure that the condition $X=\Lambda\alpha$ implies that $X$ is in fact free over $\Lambda$ (see [BHJ22, Lemma 2.3]); without this, the implication (i) $\Rightarrow$ (v) in Proposition 7.1 may not hold. This omission does not affect any of the other results of [BHJ22].
Corollary 7.3.
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Then the following are equivalent:
  1. (i)
    $X$ is free over $\Lambda$,
  2. (ii)
    there exists $u\in\mathcal{M}^{\times}$ and $\overline{\alpha}\in(\Lambda/\mathfrak{f})^{\times}$ such that $\bar{u}=\bar{\beta}\bar{\alpha}$,
  3. (iii)
    there exists $u\in\mathcal{M}^{\times}$ such that $\pi(\bar{\beta})=\pi(\bar{u})$.
Proof.
This follows from Proposition 7.1 after substituting $u$ and $\bar{\alpha}$ for their inverses. ∎
7.3. An exact sequence
Most of the rest of this section shall be concerned with the development of a criterion for stable freeness. We shall first need a certain exact sequence.
Let
$\mathcal{D}=\mathcal{D}(\Lambda,\mathfrak{f}):=\{(\lambda_{1},\lambda_{2})\in\Lambda\oplus\Lambda\mid\lambda_{1}\equiv\lambda_{2}\bmod{\mathfrak{f}}\}.$
Let $g_{1}=g_{2}:\Lambda\rightarrow\Lambda/\mathfrak{f}$ be the canonical projection map. For $i=1,2$ define $\pi_{i}:\mathcal{D}\rightarrow\Lambda$ to be the restriction of the projection map from $\Lambda\oplus\Lambda$ onto its $i$th component. Then we have a fibre-product
${\mathcal{D}}$${\Lambda}$${\Lambda}$${\Lambda/\mathfrak{f}.}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{g_{1}}$$\scriptstyle{g_{2}}$ (7.2)
For $i=1,2$ define $q_{i}:\mathrm{Cl}(\mathcal{D})\rightarrow\mathrm{Cl}(\Lambda)$ by $[X]_{\mathrm{st}}\mapsto[\Lambda\otimes_{\mathcal{D},\pi_{i}}X]_{\mathrm{st}}$. Then by Proposition 5.6 we obtain an exact sequence
$1\longrightarrow\ker(\partial)\longrightarrow(\Lambda/\mathfrak{f})^{\times}\stackrel{{\scriptstyle\partial}}{{\longrightarrow}}\mathrm{Cl}(\mathcal{D})\xrightarrow{(q_{1},q_{2})}\mathrm{Cl}(\Lambda)\oplus\mathrm{Cl}(\Lambda)\longrightarrow 0,$
where $\partial(u)=[M(u)]_{\mathrm{st}}$ and $M(u):=\{(\lambda_{1},\lambda_{2})\in\Lambda\oplus\Lambda\mid g_{1}(\lambda_{1})u=g_{2}(\lambda_{2})\}$. Set
$\mathrm{Cl}(\Lambda,\mathfrak{f}):=\ker(q_{2}:\mathrm{Cl}(\mathcal{D})\rightarrow\mathrm{Cl}(\Lambda)).$
Then we obtain an exact sequence
$(\Lambda/\mathfrak{f})^{\times}\stackrel{{\scriptstyle\partial}}{{\longrightarrow}}\mathrm{Cl}(\Lambda,\mathfrak{f})\stackrel{{\scriptstyle q_{1}}}{{\longrightarrow}}\mathrm{Cl}(\Lambda)\longrightarrow 0.$ (7.3)
7.4. Decomposition via central idempotents
Much of the following notation is adopted from [BHJ22]. Denote the centre of a ring $R$ by $Z(R)$. Let $C=Z(A)$ and let $\mathcal{O}_{C}$ be the integral closure of $\mathcal{O}$ in $C$. Set $\mathfrak{g}:=\mathfrak{f}\cap\mathcal{O}_{C}$ and note that this is a proper full ideal of $\mathcal{O}_{C}$. Let $e_{1},\ldots,e_{r}$ be the primitive idempotents of $C$ and set $A_{i}=Ae_{i}$. Then
$A=A_{1}\oplus\cdots\oplus A_{r}$ (7.4)
is a decomposition of $A$ into indecomposable two-sided ideals (see [CR81, (3.22)]). Each $A_{i}$ is a simple $K$-algebra with identity element $e_{i}$. The centres $K_{i}:=Z(A_{i})$ are finite field extensions of $K$ via $K\rightarrow K_{i}$, $\alpha\mapsto\alpha e_{i}$, and we have $K$-algebra isomorphisms $A_{i}\cong\mathrm{Mat}_{n_{i}}(D_{i})$ where $D_{i}$ is a skew field with $Z(D_{i})\cong K_{i}$ (see [CR81, (3.28)]). The Wedderburn decomposition (7.4) induces decompositions
$C=K_{1}\oplus\cdots\oplus K_{r}\quad\textrm{and}\quad\mathcal{O}_{C}=\mathcal{O}_{K_{1}}\oplus\cdots\oplus\mathcal{O}_{K_{r}},$ (7.5)
where $\mathcal{O}_{K_{i}}$ denotes the ring of algebraic integers of $K_{i}$. By [Rei03, (10.5)] we have $e_{1},\ldots,e_{r}\in\mathcal{M}$ and each $\mathcal{M}_{i}:=\mathcal{M}e_{i}$ is a maximal $\mathcal{O}$-order (and thus a maximal $\mathcal{O}_{K_{i}}$-order) in $A_{i}$. Moreover, each $\mathfrak{f}_{i}:=\mathfrak{f}e_{i}$ is a full two-sided ideal of $\mathcal{M}_{i}$, each $\mathfrak{g}_{i}:=\mathfrak{g}e_{i}$ is a nonzero integral ideal of $\mathcal{O}_{K_{i}}$, and we have decompositions
$\mathcal{M}=\mathcal{M}_{1}\oplus\cdots\oplus\mathcal{M}_{r},\quad\mathfrak{f}=\mathfrak{f}_{1}\oplus\cdots\oplus\mathfrak{f}_{r}\quad\textrm{and}\quad\mathfrak{g}=\mathfrak{g}_{1}\oplus\cdots\oplus\mathfrak{g}_{r}.$ (7.6)
7.5. The reduced norm map
The reduced norm map $\mathrm{nr}:A\rightarrow C$ is defined componentwise (see [Rei03, §9]) and restricts to a group homomorphism $\mathrm{nr}:A^{\times}\rightarrow C^{\times}$. By the Hasse-Schilling-Maas Theorem [CR81, Theorem 7.48] we have $\mathrm{nr}(A^{\times})=C^{\times+}$, where
$C^{\times+}:=\{c\in C^{\times}\mid c\text{ is positive at quaternionic components}\}.$
The last condition means that if $c=(c_{i})$ with $c_{i}\in K_{i}^{\times}$ then $\tau(c_{i})>0$ whenever $i\in\{1,\ldots,r\}$ and $\tau\colon K_{i}\to\mathbb{R}$ is a real embedding such that the corresponding scalar extension $A_{i}\otimes_{K_{i},\tau}\mathbb{R}$ is a full matrix ring over the quaternions. By [CR87, (45.7)] we have that $\mathrm{nr}(\mathcal{M}^{\times})=\mathcal{O}_{C}^{\times}\cap C^{\times+}$.
7.6. Idèles
Let $\mathfrak{p}$ range over all maximal ideals of $\mathcal{O}$. Let $K_{\mathfrak{p}}$ denote the $\mathfrak{p}$-adic completion of $K$ and let $A_{\mathfrak{p}}=K_{\mathfrak{p}}\otimes_{K}A$. Recall from §2.2 that for an $\mathcal{O}$-lattice $M$, we write $M_{\mathfrak{p}}$ and $\widehat{M}_{\mathfrak{p}}$ for the localisation and completion of $M$ at $\mathfrak{p}$, respectively.
The group of idèles of $A$ is defined to be
$J(A)=\{(a_{\mathfrak{p}})_{\mathfrak{p}}\in\textstyle{\prod_{\mathfrak{p}}}A^{\times}_{\mathfrak{p}}\mid a_{\mathfrak{p}}\in\widehat{\Lambda}^{\times}_{\mathfrak{p}}\text{ for almost all }\mathfrak{p}\}.$
Note that $J(A)$ does not depend on the choice of $\mathcal{O}$-order $\Lambda$ in $A$. Let $U(\Lambda)$ be the subgroup of unit idèles, defined by $U(\Lambda)=\prod_{\mathfrak{p}}\widehat{\Lambda}^{\times}_{\mathfrak{p}}$. Let $U_{\mathfrak{f}}(\Lambda)$ be the subgroup of $U(\Lambda)$ defined by
$U_{\mathfrak{f}}(\Lambda):=\{(a_{\mathfrak{p}})_{\mathfrak{p}}\in U(\Lambda)\mid a_{\mathfrak{p}}\equiv 1\bmod\mathfrak{f}_{\mathfrak{p}}\}.$
One has canonical isomorphisms
$A_{\mathfrak{p}}=K_{\mathfrak{p}}\otimes_{K}A$ $\cong\bigoplus_{i=1}^{r}K_{\mathfrak{p}}\otimes_{K}A_{i}\cong\bigoplus_{i=1}^{r}K_{\mathfrak{p}}\otimes_{K}K_{i}\otimes_{K_{i}}A_{i}$
$\cong\bigoplus_{i=1}^{r}\bigoplus_{\mathfrak{P}}\ (K_{i})_{\mathfrak{P}}\otimes_{K_{i}}A_{i}\cong\bigoplus_{i,\mathfrak{P}}A_{i,\mathfrak{P}}$ (7.7)
involving various completions, where, for given $i\in\{1,\ldots,r\}$, $\mathfrak{P}$ runs through all maximal ideals of $\mathcal{O}_{K_{i}}$ dividing $\mathfrak{p}$ and $A_{i,\mathfrak{P}}$ is defined as $(A_{i})_{\mathfrak{P}}$. Using the above isomorphism, we will often write elements of $J(A)$, resp. $A_{\mathfrak{p}}$, as tuples $(a_{i,\mathfrak{P}})_{i,\mathfrak{P}}$, where $\mathfrak{P}$ ranges over all maximal ideals of $\mathcal{O}_{K_{i}}$, resp. over those that contain $\mathfrak{p}$. Similarly, we denote by $J(C)$ the group of idèles of $C$. Again one has a canonical isomorphism
$C_{\mathfrak{p}}\cong\bigoplus_{i,\mathfrak{P}}\ K_{i,\mathfrak{P}}$
and we will write elements in $J(C)$, resp. $C_{\mathfrak{p}}$, often in the form $(\alpha_{i,\mathfrak{P}})_{i,\mathfrak{P}}$.
By $\mathrm{nr}\colon J(A)\to J(C)$ we also denote the reduced norm map (which translates into the component-wise reduced norm maps $\mathrm{nr}\colon A_{i,\mathfrak{P}}^{\times}\to K_{i,\mathfrak{P}}^{\times}$ after the above identifications). Then by [CR87, (45.8)] we have that $\mathrm{nr}(J(A))=J(C)$.
Theorem 7.4 ([CR87, (49.17)]).
There is a canonical isomorphism
$\mathrm{Cl}(\Lambda)\cong\frac{J(C)}{C^{\times+}\mathrm{nr}(U(\Lambda))}.$
Remark 7.5.
For $\gamma=(\gamma_{\mathfrak{p}})_{\mathfrak{p}}\in J(A)$ define $\Lambda\gamma=A\cap(\cap_{\mathfrak{p}}\widehat{\Lambda}_{\mathfrak{p}}\gamma_{\mathfrak{p}})$. Then the map $\varepsilon:J(A)\rightarrow\mathrm{Cl}(\Lambda)$ defined by $\alpha\mapsto[\Lambda\alpha]_{\mathrm{st}}$ is a surjective homomorphism. Now let $X$ be a locally free left ideal of $\Lambda$. Then for each $\mathfrak{p}$ there exists $\alpha_{\mathfrak{p}}\in A_{\mathfrak{p}}^{\times}\cap\widehat{\Lambda}_{\mathfrak{p}}$ such $\widehat{X}_{\mathfrak{p}}=\widehat{\Lambda}_{\mathfrak{p}}\alpha_{\mathfrak{p}}$. Set $\alpha=(\alpha_{\mathfrak{p}})_{\mathfrak{p}}$. Then $\Lambda\alpha=X$ and so $\varepsilon(\alpha)=[X]_{\mathrm{st}}$. In fact, $\varepsilon$ factors via $\mathrm{nr}\colon J(A)\to J(C)$ and so under the isomorphism of Theorem 7.4 $(\mathrm{nr}(\alpha_{\mathfrak{p}})_{\mathfrak{p}})\in J(C)$ is a representative of $[X]_{\mathrm{st}}$. Note that we could also have chosen $\alpha_{\mathfrak{p}}\in A^{\times}\cap\Lambda_{\mathfrak{p}}$ such that $X_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}\alpha_{\mathfrak{p}}$. For further details on the claims in this remark, see [CR87, §49A].
7.7. Review of results of Bley-Boltje
We now recall some key results from [BB06].
Let $I_{\mathfrak{g}}=I_{\mathfrak{g}}(C)$ denote the group of fractional $\mathcal{O}_{C}$-ideals of $C$ that are coprime to $\mathfrak{g}$, and for each $i$, define $I_{\mathfrak{g}_{i}}(K_{i})$ analogously. Then we have a direct product decomposition
$I_{\mathfrak{g}}(C)=I_{\mathfrak{g}_{1}}(K_{1})\times\cdots\times I_{\mathfrak{g}_{r}}(K_{r}).$
For each $i\in\{1,\ldots,r\}$ we write $\infty_{i}$ for the formal product over real archimedean places $\tau\colon K_{i}\to\mathbb{R}$ such that $\mathbb{R}\otimes_{K_{i},\tau}A_{i}$ is a full matrix ring over the quaternions, and we define the ‘ray modulo $\mathfrak{g}\infty$’ by
$P_{\mathfrak{g}}^{+}=P_{\mathfrak{g}}^{+}(C)=\{(\alpha_{i}\mathcal{O}_{K_{i}})_{i}\in I_{\mathfrak{g}}\mid\alpha_{i}\equiv 1\bmod\!\!^{\times}\mathfrak{g}_{i}\infty_{i}\text{ for all }i=1,\ldots,r\}.$
Then $P_{\mathfrak{g}}^{+}$ is a subgroup of $I_{\mathfrak{g}}$ and we have a corresponding direct product decomposition
$P_{\mathfrak{g}}(C)^{+}=P_{\mathfrak{g}_{1}}^{+}(K_{1})\times\cdots\times P_{\mathfrak{g}_{r}}^{+}(K_{r}).$
For each $i$, the quotient $I_{\mathfrak{g}_{i}}(K_{i})/P_{\mathfrak{g}_{i}}^{+}(K_{i})$ is a ray class group of $K_{i}$, and $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$ is the direct product of these groups.
Theorem 7.6 ([BB06, Theorem 1.5]).
There are canonical isomorphisms
$\mathrm{Cl}(\Lambda,\mathfrak{f})\cong\frac{J(C)}{C^{\times+}\mathrm{nr}(U_{\mathfrak{f}}(\Lambda))}\cong I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}.$
Using these canonical isomorphisms, the exact sequence (7.3) can be rewritten as
$\left(\Lambda/\mathfrak{f}\right)^{\times}\stackrel{{\scriptstyle\nu}}{{\longrightarrow}}I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}\stackrel{{\scriptstyle\mu}}{{\longrightarrow}}\mathrm{Cl}(\Lambda)\longrightarrow 0.$ (7.8)
Proposition 7.7 ([BB06, Proposition 1.8]).
Let $x\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ and let $a\in\Lambda\cap A^{\times}$ such that $a\equiv x\pmod{\mathfrak{f}}$. Then $\nu(x)=(\mathrm{nr}(a)\mathcal{O}_{C}\bmod P_{\mathfrak{g}}^{+})$ in $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$.
Let $x\in\left(\mathcal{M}/\mathfrak{f}\right)^{\times}$ and let $a\in\mathcal{M}\cap A^{\times}$ such that $a\equiv x\pmod{\mathfrak{f}}$. Then define
$\tilde{\nu}(x)=(\mathrm{nr}(a)\mathcal{O}_{C}\bmod P_{\mathfrak{g}}^{+})\in I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}.$ (7.9)
Thus we obtain a group homomorphism $\tilde{\nu}:(\mathcal{M}/\mathfrak{f})^{\times}\longrightarrow I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$. By Proposition 7.7, $\tilde{\nu}$ extends $\nu$, that is, $\tilde{\nu}(x)=\nu(x)$ for $x\in(\Lambda/\mathfrak{f})^{\times}$. Note that
$\ker(\mu)=\nu((\Lambda/\mathfrak{f})^{\times})=\tilde{\nu}((\Lambda/\mathfrak{f})^{\times}).$ (7.10)
7.8. A condition on $\mathfrak{g}$
For a maximal ideal $\mathfrak{P}$ of $\mathcal{O}_{C}=\mathcal{O}_{K_{1}}\oplus\dotsb\oplus\mathcal{O}_{K_{r}}$ and a finitely generated $\mathcal{O}_{C}$-module $M$ we write $M_{\mathfrak{P}}$ for the localisation of $M$ at $\mathfrak{P}$. Note that for a given $\mathfrak{P}$ there exists a unique $1\leq i\leq r$ and a maximal $\mathcal{O}_{K_{i}}$-ideal $\mathfrak{P}_{i}$ such that
$\mathfrak{P}=\mathcal{O}_{K_{1}}\oplus\dotsb\oplus\mathcal{O}_{K_{i-1}}\oplus\mathfrak{P}_{i}\oplus\mathcal{O}_{K_{i+1}}\oplus\dotsb\oplus\mathcal{O}_{K_{r}}.$
We can and will henceforth identify $M_{\mathfrak{P}}$ with the localisation $(e_{i}M)_{\mathfrak{P}_{i}}$ of the $\mathcal{O}_{K_{i}}$-module $e_{i}M$ at $\mathfrak{P}_{i}$. Note that $M$ is uniquely determined by $M_{\mathfrak{P}}$ for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{C}$ and $M_{\mathfrak{p}}$, where $\mathfrak{p}$ is a maximal ideal of $\mathcal{O}$, is uniquely determined by $M_{\mathfrak{P}}$ for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{C}$ lying above $\mathfrak{p}$.
We shall need the following assumption on $\mathfrak{g}$ in the proof of Theorem 7.10 below.
Assumption (G).
If $\mathfrak{p}$ is a maximal ideal of $\mathcal{O}$ with $\mathfrak{g}_{\mathfrak{p}}\neq\mathcal{O}_{C,\mathfrak{p}}$, then we have $\mathfrak{g}_{\mathfrak{P}}\subseteq\mathfrak{P}\mathcal{O}_{C,\mathfrak{P}}$ for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{C}$ with $\mathfrak{P}\mid\mathfrak{p}$.
Example 7.8.
Let $L/K$ be an extension of number fields, let $\mathfrak{p}$ be a maximal ideal of $\mathcal{O}_{K}$ and let $\mathfrak{P}$ be a prime ideal of $\mathcal{O}_{L}$ lying over $\mathfrak{p}$. We consider the $\mathcal{O}_{K}$-order $\Lambda=\mathcal{O}_{K}+\mathfrak{P}$ and the (unique) maximal $\mathcal{O}_{K}$-order $\mathcal{M}=\mathcal{O}_{L}$ in the $K$-algebra $L$. Then $A=C=L$ and for $\mathfrak{g}=\mathfrak{f}=\mathfrak{P}$ we have $\mathfrak{f}_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}=\mathcal{O}_{C,\mathfrak{p}}$ and assumption (G) is not satisfied whenever there are at least two prime ideals of $\mathcal{O}_{L}$ lying over $\mathfrak{p}$.
Lemma 7.9.
The following statements hold.
  1. (i)
    Let $\mathfrak{p}$ be a maximal ideal of $\mathcal{O}$. Then $\mathfrak{f}_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}$ if and only if $\mathfrak{g}_{\mathfrak{p}}=\mathcal{O}_{C,\mathfrak{p}}$.
  2. (ii)
    Let $\mathfrak{P}$ be a maximal ideal of $\mathcal{O}_{C}$. Then $\mathfrak{f}_{\mathfrak{P}}\subseteq\mathrm{rad}(\mathcal{M}_{\mathfrak{P}})$ if and only if $\mathfrak{g}_{\mathfrak{P}}\subseteq\mathfrak{P}\mathcal{O}_{C,\mathfrak{P}}$.
Proof.
We have $\mathfrak{g}_{\mathfrak{p}}=(\mathfrak{f}\cap\mathcal{O}_{C})_{\mathfrak{p}}=\mathfrak{f}_{\mathfrak{p}}\cap\mathcal{O}_{C,\mathfrak{p}}$ since localisations commute with intersections. Thus $1\in\mathfrak{g}_{\mathfrak{p}}$ if and only if $1\in\mathfrak{f}_{\mathfrak{p}}$, which immediately implies (i). Substituting $\mathfrak{P}$ for $\mathfrak{p}$ in this argument shows that $\mathfrak{f}_{\mathfrak{P}}=\mathcal{M}_{\mathfrak{P}}$ if and only if $\mathfrak{g}_{\mathfrak{P}}=\mathcal{O}_{C,\mathfrak{P}}$. By [Rei03, (18.3)], we have that $\mathrm{rad}(\mathcal{M}_{\mathfrak{P}})$ is the unique maximal two-sided ideal of $\mathcal{M}_{\mathfrak{P}}$. Moreover, $\mathfrak{P}$ is clearly the unique maximal ideal of $\mathcal{O}_{C,\mathfrak{P}}$. Together, these last three facts imply (ii). ∎
7.9. A criterion for stable freeness
Recall that the maps $\mu$ and $\tilde{\nu}$ were defined in (7.8) and (7.9), respectively, and that for $\eta\in\mathcal{M}$ we write $\overline{\eta}$ for its image in $\mathcal{M}/\mathfrak{f}$.
Theorem 7.10.
Assume that $\mathfrak{g}$ satisfies Assumption (G). Let $X$ be a locally free left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Then
$\mu(\tilde{\nu}(\overline{\beta}))=\mu(\mathrm{nr}(\beta)\mathcal{O}_{C}\bmod P_{\mathfrak{g}}^{+})=[X]_{\mathrm{st}}\text{ in }\mathrm{Cl}(\Lambda).$
Proof.
Let $\mathfrak{p}$ range over all maximal ideals of $\mathcal{O}$. For each $\mathfrak{p}$, we define an element $\alpha_{\mathfrak{p}}\in X_{\mathfrak{p}}$ with $X_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}\alpha_{\mathfrak{p}}$ as follows. If $\mathfrak{f}_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}$ then we set $\alpha_{\mathfrak{p}}:=\beta$. If $\mathfrak{f}_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$, we let $\alpha_{\mathfrak{p}}^{\prime}\in X_{\mathfrak{p}}$ be an arbitrarily chosen element such that $X_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}\alpha_{\mathfrak{p}}^{\prime}$. Then $\Lambda_{\mathfrak{p}}\alpha_{\mathfrak{p}}^{\prime}+\mathfrak{f}_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}$, so $\overline{\alpha}_{\mathfrak{p}}^{\prime}\in\left(\Lambda_{\mathfrak{p}}/\mathfrak{f}_{\mathfrak{p}}\right)^{\times}$. Since $\Lambda_{\mathfrak{p}}$ is semilocal, the canonical map $\Lambda_{\mathfrak{p}}^{\times}\longrightarrow(\Lambda_{\mathfrak{p}}/\mathfrak{f}_{\mathfrak{p}})^{\times}$ is surjective by [Bas68, Chapter III, (2.9) Corollary]. In particular, there exists $\alpha_{\mathfrak{p}}^{\prime\prime}\in\Lambda_{\mathfrak{p}}^{\times}$ such that $\alpha_{\mathfrak{p}}^{\prime}\equiv\alpha_{\mathfrak{p}}^{\prime\prime}\pmod{\mathfrak{f}_{\mathfrak{p}}}$. Then we set $\alpha_{\mathfrak{p}}:=(\alpha_{\mathfrak{p}}^{\prime\prime})^{-1}\alpha_{\mathfrak{p}}^{\prime}$ and note that $\alpha_{\mathfrak{p}}\equiv 1\pmod{\mathfrak{f}_{\mathfrak{p}}}$.
We have the following commutative diagram
[…]
where the arrows in the upper triangle are the canonical projections, the isomorphism on the right is that of Theorem 7.4, and the two isomorphisms on the left are those of Theorem 7.6, and $\psi$ is defined to be their composite. Recall that $\mathrm{nr}$, $q_{1}$, $\mu$, and $\varepsilon$ are defined in §7.6, (7.3), (7.8), and Remark 7.5, respectively. The only part of the diagram for which commutativity is not immediately implied by the definitions and Remark 7.5 is the central rectangle, but this follows from the definitions of the isomorphisms as discussed in the proof of [BB06, Theorem 1.5].
By Remark 7.5 and the commutativity of the diagram we see that
$\mu\circ\psi\left((\mathrm{nr}(\alpha_{\mathfrak{p}}))_{\mathfrak{p}}\bmod C^{\times+}\mathrm{nr}(U_{\mathfrak{f}}(\Lambda))\right)=[X]_{\mathrm{st}}.$
We recall the recipe for the computation of $\psi$ from the proof of [BB06, Theorem 1.5]. By [BB06, Remark 1.6] we have $U_{\mathfrak{f}}(\Lambda)=U_{\mathfrak{f}}(\mathcal{M})$. Hence both the domain and codomain of $\psi$ break up into direct sums of components, one for each $i=1,\ldots,r$. So we fix $i\in\{1,\ldots,r\}$, let $L:=K_{i}$ and replace $\mathfrak{g}_{i}$ by $\mathfrak{g}$, $\mathfrak{f}_{i}$ by $\mathfrak{f}$, $\mathcal{M}_{i}$ by $\mathcal{M}$ and $A_{i}$ by $A$.
Let $\mathfrak{P}$ range over all maximal ideals of $\mathcal{O}_{L}$. Let $\gamma=\left(\gamma_{\mathfrak{P}}\right)_{\mathfrak{P}}\in J(L)$. In order to compute $\psi(\gamma\bmod L^{\times+}\mathrm{nr}(U_{\mathfrak{f}}(\mathcal{M})))$ we note that by the Weak Approximation Theorem [Coh00, Proposition 1.2.3 (3)] there exists $\xi\in L^{\times+}$ (named $\beta$ in [BB06]) such that
$v_{\mathfrak{P}}(\gamma_{\mathfrak{P}}\xi-1)\geq v_{\mathfrak{P}}(\mathfrak{g})\text{ for all }\mathfrak{P}\text{ with }\mathfrak{P}\mid\mathfrak{g}.$
Then the class of
$\mathfrak{a}=\prod_{\mathfrak{P}}\mathfrak{P}^{v_{\mathfrak{P}}(\gamma_{\mathfrak{P}}\xi)}$
is a representative of $\psi(\gamma\bmod L^{\times+}\mathrm{nr}(U_{\mathfrak{f}}(\mathcal{M})))$ in $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$. In the definition of $\mathfrak{a}$ the product extends over all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{L}$.
Using the canonical isomorphism of (10.2) we write $\alpha_{\mathfrak{p}}=\left(\alpha_{\mathfrak{P}}\right)_{\mathfrak{P}\mid\mathfrak{p}}$ with $\alpha_{\mathfrak{P}}\in A_{\mathfrak{P}}$ for the local basis elements $\alpha_{\mathfrak{p}}$ chosen as above. Then, for $\mathfrak{P}\mid\mathfrak{g}$, the result of [BB06, Corollary 2.4] implies $\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}})\equiv 1\pmod{\mathfrak{g}_{\mathfrak{P}}}$, so that we can and do choose $\xi=1$.
It suffices to show that
${v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}}))}=v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\beta))$ (7.11)
for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{L}$. To this end, fix a maximal ideal $\mathfrak{p}$ of $\mathcal{O}$.
If $\mathfrak{f}_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}$, then $\alpha_{\mathfrak{P}}=\beta$ for all $\mathfrak{P}\mid\mathfrak{p}$ and (7.11) is clear. Now suppose $\mathfrak{f}_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$. First note that we have
$\mathcal{M}_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}X_{\mathfrak{p}}+\mathfrak{f}_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}\beta+\mathfrak{f}_{\mathfrak{p}}.$
Hence there exists an element $\gamma\in\mathcal{M}_{\mathfrak{p}}$ such that $\gamma\beta\equiv 1\pmod{\mathfrak{f}_{\mathfrak{p}}}$. It follows that $\gamma\beta\equiv 1\pmod{\mathfrak{f}_{\mathfrak{P}}}$ for all $\mathfrak{P}\mid\mathfrak{p}$. From [BB06, Corollary 2.4] we obtain $\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\gamma\beta)\equiv 1\pmod{\mathfrak{g}_{\mathfrak{P}}}$ for all $\mathfrak{P}\mid\mathfrak{p}$. Since $\mathfrak{g}_{\mathfrak{p}}\neq\mathcal{O}_{C,\mathfrak{p}}$ by Lemma 7.9, assumption (G) implies that $\mathfrak{g}_{\mathfrak{P}}\subseteq\mathfrak{P}\mathcal{O}_{C,\mathfrak{P}}$. Hence $v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\beta))=0$. Since $\alpha_{\mathfrak{p}}\equiv 1\pmod{\mathfrak{f}_{\mathfrak{p}}}$ the same argument shows that $v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}}))=0$ for all maximal ideals $\mathfrak{P}\mid\mathfrak{p}$.
Hence (7.11) holds for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{L}$. ∎
7.10. Choosing $\mathfrak{g}$ first
Recall that in §7.1 and §7.4, respectively, we first chose $\mathfrak{f}$ to be a proper full two-sided ideal of $\mathcal{M}$ that is contained in $\Lambda$ and then set $\mathfrak{g}:=\mathfrak{f}\cap\mathcal{O}_{C}$. In some applications, it is convenient to work the other way round: first choose an proper full ideal $\mathfrak{g}$ of $\mathcal{O}_{C}$ such that $\mathfrak{g}\mathcal{M}\subseteq\Lambda$ and then set $\mathfrak{f}:=\mathfrak{g}\mathcal{M}$. The following lemma shows that the latter approach is consistent with the former.
Lemma 7.11.
Let $\mathfrak{g}$ be a proper full ideal of $\mathcal{O}_{C}$ such that $\mathfrak{f}:=\mathfrak{g}\mathcal{M}\subseteq\Lambda$. Then $\mathfrak{f}$ is a proper full two sided-ideal of $\mathcal{M}$ and $\mathfrak{f}\cap\mathcal{O}_{C}=\mathfrak{g}$.
Proof.
The first claim is clear. To prove the second claim, we can and do work componentwise and hence assume without loss of generality that $A=\mathrm{Mat}_{n}(D)$ for some $n\geq 1$ and some skewfield $D$ containing $K$. Let $F=Z(D)$ and note that $F=C$.
Let $\Delta$ be any maximal $\mathcal{O}_{F}$-order in $D$. Then $\Delta\cap F=\mathcal{O}_{F}$. Thus if $\alpha\delta\in\mathcal{O}_{F}$ for some $\delta\in\Delta$ and some $\alpha\in\mathcal{O}_{F}$ with $\alpha\neq 0$ then $\delta=\alpha^{-1}(\alpha\delta)\in\Delta\cap F=\mathcal{O}_{F}$. This implies that as an $\mathcal{O}_{F}$-module, $\Delta/\mathcal{O}_{F}$ is torsion-free and hence is projective. Thus the canonical short exact sequence of $\mathcal{O}_{F}$-modules
$0\longrightarrow\mathcal{O}_{F}\longrightarrow\Delta\longrightarrow\Delta/\mathcal{O}_{F}\longrightarrow 0$
splits, which gives an isomorphism of $\mathcal{O}_{F}$-lattices $\Delta\cong\mathcal{O}_{F}\oplus\Delta/\mathcal{O}_{F}$. Abusing notation slightly, it follows that
$\mathfrak{g}\Delta\cap\mathcal{O}_{F}=\mathfrak{g}\mathcal{O}_{F}\oplus\mathfrak{g}(\Delta/\mathcal{O}_{F})\cap\mathcal{O}_{F}=\mathfrak{g}.$ (7.12)
If $n=1$ then $\mathcal{M}$ is a maximal $\mathcal{O}_{F}$-order in $D=A$ and so taking $\Delta=\mathcal{M}$ in equation (7.12) gives the desired result.
Suppose that $n\geq 2$ and again let $\Delta$ be any maximal $\mathcal{O}_{F}$-order in $D$. For each right ideal $\mathfrak{a}$ of $\Delta$, let $\Delta^{\prime}=\mathcal{O}_{l}(\mathfrak{a}):=\{x\in D\mid x\mathfrak{a}\subseteq\mathfrak{a}\}$, and let
$\mathcal{M}_{\mathfrak{a},n}=\left(\begin{array}[]{cccc}\Delta&\ldots&\Delta&\mathfrak{a}^{-1}\\ \vdots&\ddots&\vdots&\vdots\\ \Delta&\ldots&\Delta&\mathfrak{a}^{-1}\\ \mathfrak{a}&\ldots&\mathfrak{a}&\Delta^{\prime}\end{array}\right)$
denote the ring of all $n\times n$ matrices $(x_{ij})$ where $x_{11}$ ranges over all elements of $\Delta$, …, $x_{1n}$ ranges over all elements of $\mathfrak{a}^{-1}$, and so on. By [Rei03, (27.6)], $\mathcal{M}_{\mathfrak{a},n}$ is a maximal $\mathcal{O}_{F}$-order in $A$ and there exists a choice of $\mathfrak{a}$ such that there is an isomorphism of $\mathcal{O}_{F}$-orders $\theta:\mathcal{M}_{\mathfrak{a},n}\cong\mathcal{M}$. Since $\Delta$ was arbitrary in (7.12), we have $\mathfrak{g}\Delta\cap\mathcal{O}_{F}=\mathfrak{g}=\mathfrak{g}\Delta^{\prime}\cap\mathcal{O}_{F}$. Thus it follows easily from the explicit description of $\mathcal{M}_{\mathfrak{a},n}$ that we have $\mathfrak{g}\mathcal{M}_{\mathfrak{a},n}\cap\mathcal{O}_{F}=\mathfrak{g}$. Since $\theta(\mathcal{O}_{F})=\mathcal{O}_{F}$, $\theta(\mathfrak{g})=\mathfrak{g}$ and $\theta(\mathfrak{g}\mathcal{M}_{\mathfrak{a},n})=\mathfrak{g}\mathcal{M}$, we conclude that $\mathfrak{g}\mathcal{M}\cap\mathcal{O}_{F}=\mathfrak{g}$, as desired. ∎
8. An algorithm for determining SFC
8.1. Generalised Swan modules
We assume the setup and notation of §7.
Definition 8.1.
For $\beta\in\mathcal{M}\cap A^{\times}$ we define $X_{\beta}=X_{{\beta},\Lambda,\mathcal{M}}=\mathcal{M}\beta\cap\Lambda$, and call this the generalised Swan module attached to $\beta$ (and $\Lambda,\mathcal{M}$). Note that $X_{\beta}$ is a left ideal of $\Lambda$.
The following lemma determines the localisations of certain generalised Swan modules.
Lemma 8.2.
Suppose that $\mathfrak{g}$ satisfies Assumption (G). Let $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}\beta+\mathfrak{f}=\mathcal{M}$. Let $\mathfrak{p}$ be a maximal ideal of $\mathcal{O}$. Then
$(X_{\beta})_{\mathfrak{p}}=\begin{cases}\Lambda_{\mathfrak{p}}\beta&\text{ if }\Lambda_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}},\\ \Lambda_{\mathfrak{p}}&\text{ if }\Lambda_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}.\end{cases}$
Moreover, if $\Lambda_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$ then $\beta\in\mathcal{M}_{\mathfrak{p}}^{\times}$.
Proof.
If $\Lambda_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}$ then $(X_{\beta})_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}\beta\cap\Lambda_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}\beta\cap\mathcal{M}_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}\beta=\Lambda_{\mathfrak{p}}\beta$.
Now suppose that $\Lambda_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$. Then $\mathfrak{f}_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$ and so by Assumption (G) and Lemma 7.9 and we have that $\mathfrak{f}_{\mathfrak{P}}\subseteq\mathrm{rad}(\mathcal{M}_{\mathfrak{P}})$ for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{C}$ with $\mathfrak{P}\mid\mathfrak{p}$. Then $\mathcal{M}_{\mathfrak{P}}\beta+\mathfrak{f}_{\mathfrak{P}}=(\mathcal{M}+\mathfrak{f})_{\mathfrak{P}}=\mathcal{M}_{\mathfrak{P}}$ and so Nakayama’s lemma implies that $\mathcal{M}_{\mathfrak{P}}\beta=\mathcal{M}_{\mathfrak{P}}$ for all $\mathfrak{P}\mid\mathfrak{p}$, or equivalently, $\beta\in\mathcal{M}_{\mathfrak{p}}^{\times}$. Hence $\mathcal{M}_{\mathfrak{p}}\beta=\mathcal{M}_{\mathfrak{p}}$ and so
$(X_{\beta})_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}\beta\cap\Lambda_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}\cap\Lambda_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}.\qed$
The following result establishes further properties of certain generalised Swan modules. We recall that the maps $\pi,\mu$ and $\tilde{\nu}$ were defined in (7.1), (7.8) and (7.9), respectively.
Proposition 8.3.
Suppose that $\mathfrak{g}$ satisfies Assumption (G). Let $\overline{\beta}\in\left(\mathcal{M}/\mathfrak{f}\right)^{\times}$. Then for every lift $\beta\in\mathcal{M}\cap A^{\times}$ we have that
  1. (i)
    $\mathcal{M}\beta+\mathfrak{f}=\mathcal{M}$,
  2. (ii)
    $X_{\beta}$ is locally free over $\Lambda$,
  3. (iii)
    $\mathcal{M}X_{\beta}=\mathcal{M}\beta$,
  4. (iv)
    $X_{\beta}+\mathfrak{f}=\Lambda$,
  5. (v)
    $\tilde{\nu}(\overline{\beta})=(\mathrm{nr}(\beta)\mathcal{O}_{C}\bmod P_{\mathfrak{g}}^{+})$ in $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$,
  6. (vi)
    $\mu(\tilde{\nu}(\overline{\beta}))=[X_{\beta}]_{\mathrm{st}}$ in $\mathrm{Cl}(\Lambda)$,
  7. (vii)
    $X_{\beta}$ is stably free over $\Lambda$ if and only if $\tilde{\nu}(\overline{\beta})\in\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})=\ker(\mu)$, and
  8. (viii)
    $X_{\beta}$ is free over $\Lambda$ if and only if there exists $u\in\mathcal{M}^{\times}$ such that $\pi(\overline{\beta})=\pi(\overline{u})$.
Proof.
Assertion (i) is equivalent to the hypothesis that $\overline{\beta}\in\left(\mathcal{M}/\mathfrak{f}\right)^{\times}$. Assertion (ii) follows trivially from Lemma 8.2.
Let $\mathfrak{p}$ be a maximal ideal of $\mathcal{O}$. By [Rei03, Theorem 4.21], it will suffice to prove assertions (iii) and (iv) after localising at $\mathfrak{p}$. If $\Lambda_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}$ then by Lemma 8.2 we have that $(X_{\beta})_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}\beta$ and thus
$(\mathcal{M}X_{\beta})_{\mathfrak{p}}$ $=\Lambda_{\mathfrak{p}}\beta=\mathcal{M}_{\mathfrak{p}}\beta\quad\text{and}$
$(X_{\beta}+\mathfrak{f})_{\mathfrak{p}}$ $=(X_{\beta})_{\mathfrak{p}}+\mathfrak{f}_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}\beta+\mathfrak{f}_{\mathfrak{p}}=(\mathcal{M}\beta+\mathfrak{f})_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}},$
where the final equality holds by part (i). If $\Lambda_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$ then by Lemma 8.2 we have that $(X_{\beta})_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}$ and $\beta\in\mathcal{M}_{\mathfrak{p}}^{\times}$, and thus
$(\mathcal{M}X_{\beta})_{\mathfrak{p}}$ $=\mathcal{M}_{\mathfrak{p}}(X_{\beta})_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}\Lambda_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}\beta\quad\text{and}$
$(X_{\beta}+\mathfrak{f})_{\mathfrak{p}}$ $=(X_{\beta})_{\mathfrak{p}}+\mathfrak{f}_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}+\mathfrak{f}_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}.$
Asserion (v) is just the definition of $\tilde{\nu}$ (7.9). Assertion (vi) follows from (ii), (iii) and (iv) together with Theorem 7.10. We have that $X_{\beta}$ is stably free over $\Lambda$ if and only if $[X_{\beta}]_{\mathrm{st}}=0$ and by (7.10) we have $\ker(\mu)=\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$. Therefore assertion (vii) now follows from (vi). Finally, assertion (viii) follows from (iii) and (iv) together with Corollary 7.3. ∎
Corollary 8.4.
Suppose that $\mathfrak{g}$ satisfies Assumption (G). Let $X$ be a locally free left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Then $\overline{\beta}\in(\mathcal{M}/\mathfrak{f})^{\times}$ and $[X_{\beta}]_{\mathrm{st}}=[X]_{\mathrm{st}}$ in $\mathrm{Cl}(\Lambda)$.
Proof.
We have $\mathcal{M}=\mathcal{M}(X+\mathfrak{f})=\mathcal{M}X+\mathfrak{f}=\mathcal{M}\beta+\mathfrak{f}$, and so $\overline{\beta}\in(\mathcal{M}/\mathfrak{f})^{\times}$. Then by Theorem 7.10 and Proposition 8.3 (vi) we have $[X]_{\mathrm{st}}=\mu(\tilde{\nu}(\overline{\beta}))=[X_{\beta}]_{\mathrm{st}}$ in $\mathrm{Cl}(\Lambda)$. ∎
Corollary 8.5.
Suppose that $\mathfrak{g}$ satisfies Assumption (G). Let $\beta_{1},\beta_{2}\in\mathcal{M}\cap A^{\times}$ such that $\overline{\beta_{1}},\overline{\beta_{2}}\in(\mathcal{M}/\mathfrak{f})^{\times}$. Suppose that $\overline{\beta_{1}},\overline{\beta_{2}}$ have the same image in ×\$\mathcal{M}^{\times}\backslash(\mathcal{M}/\mathfrak{f})^{\times}{}{/}{}(\Lambda/\mathfrak{f})^{\times}$. Then
  1. (i)
    $X_{\beta_{1}}$ is free over $\Lambda$ if and only if $X_{\beta_{2}}$ is free over $\Lambda$, and
  2. (ii)
    $X_{\beta_{1}}$ is stably free over $\Lambda$ if and only if $X_{\beta_{2}}$ is stably free over $\Lambda$.
Proof.
By hypothesis there exist $u\in\mathcal{M}^{\times}$ and $\overline{\alpha}\in(\Lambda/\mathfrak{f})^{\times}$ such that $\overline{\beta_{1}}=\overline{u}\overline{\beta_{2}}\overline{\alpha}$ in $(\mathcal{M}/\mathfrak{f})^{\times}$. Then Proposition 8.3 (viii) implies assertion (i). By the definition of $\tilde{\nu}$ (7.9), we have that $\tilde{\nu}(\overline{u})=(\mathrm{nr}(u)\mathcal{O}_{C}\bmod{P_{\mathfrak{g}}^{+}})$, which is trivial since $\mathrm{nr}(\mathcal{M}^{\times})\subseteq\mathcal{O}_{C}^{\times}$ (see §7.5). Thus
$\tilde{\nu}(\overline{\beta_{1}})=\tilde{\nu}(\overline{u}\overline{\beta_{2}}\overline{\alpha})=\tilde{\nu}(\overline{u})\tilde{\nu}(\overline{\beta_{2}})\tilde{\nu}(\overline{\alpha})\\ =\tilde{\nu}(\overline{\beta_{2}})\tilde{\nu}(\overline{\alpha}),$
where $\tilde{\nu}(\overline{\alpha})\in\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$, and so assertion (ii) now follows from Proposition 8.3 (vii). ∎
Lemma 8.6.
Let $Y\in g(\Lambda)$. Then there exists a locally free left ideal $X$ of $\Lambda$ such that $X+\mathfrak{f}=\Lambda$ and $X\cong Y$ as $\Lambda$-lattices.
Proof.
Since $Y\in g(\Lambda)$, by Roiter’s lemma [Rei03, (27.1)] there exists a monomorphism $\varphi~{}\colon~{}Y\hookrightarrow\Lambda$ such that $(\mathfrak{f}\cap\mathcal{O})+\mathrm{Ann}_{\mathcal{O}}(\mathrm{coker}\varphi)=\mathcal{O}$. Then we can take $X=\varphi(Y)$. ∎
Theorem 8.7.
Suppose that $\mathcal{M}$ has SFC and that $\mathfrak{g}$ satisfies Assumption (G). Let $S\subseteq\mathcal{M}\cap A^{\times}$ be a set of representatives of $(\mathcal{M}/\mathfrak{f})^{\times}$. Let $T=\{\beta\in S\colon X_{\beta}\text{ is stably free over }\Lambda\}$. Then $\Lambda$ has SFC if and only if $X_{\beta}$ is free over $\Lambda$ for every $\beta\in T$.
Proof.
Suppose that $\Lambda$ has SFC. If $\beta\in T$ then $X_{\beta}$ is stably free over $\Lambda$ and so $X_{\beta}$ is in fact free over $\Lambda$.
Recall from §3.1 that in order to show that $\Lambda$ has SFC, it suffices to show that every stably free $\Lambda$-lattice of rank $1$ is in fact free. So suppose that $X$ is a stably free $\Lambda$-lattice of rank $1$, that is, $X\in g(\Lambda)$ and $[X]_{\mathrm{st}}=0$ in $\mathrm{Cl}(\Lambda)$. By Lemma 8.6 we can and do suppose without loss of generality that $X$ is a locally free left ideal of $\Lambda$ such that $X+\mathfrak{f}=\Lambda$. Moreover, $\mathcal{M}X\cong\mathcal{M}\otimes_{\Lambda}X$ is stably free of rank $1$ over $\mathcal{M}$. Since $\mathcal{M}$ has SFC by hypothesis, there exists $\beta_{1}\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta_{1}$. Then by Corollary 8.4 we have $\overline{\beta_{1}}\in(\mathcal{M}/\mathfrak{f})^{\times}$ and $[X_{\beta_{1}}]_{\mathrm{st}}=[X]_{\mathrm{st}}=0$. Hence by definition of $S$, there exists $\beta_{2}\in S$ such that $\overline{\beta_{2}}=\overline{\beta_{1}}$. Then $[X_{\beta_{2}}]_{\mathrm{st}}=[X_{\beta_{1}}]_{\mathrm{st}}=0$ by Proposition 8.3 (vi) and so in fact $\beta_{2}\in T$. Thus $X_{\beta_{2}}$ is free over $\Lambda$ by the assumption on $T$. By Proposition 8.3 (viii) this implies that there exists $u\in\mathcal{M}^{\times}$ such that $\pi(\overline{\beta_{2}})=\pi(\overline{u})$. Since $\pi(\overline{\beta_{1}})=\pi(\overline{\beta_{2}})$, we have $\pi(\overline{\beta_{1}})=\pi(\overline{u})$. Thus $X$ is free over $\Lambda$ by Corollary 7.3. ∎
Corollary 8.8.
Suppose that $\mathcal{M}$ has SFC and that $\mathfrak{g}$ satisfies Assumption (G). Let $S\subseteq\mathcal{M}\cap A^{\times}$ be a set of representatives of ×\$\mathcal{M}^{\times}\backslash(\mathcal{M}/\mathfrak{f})^{\times}{}{/}{}(\Lambda/\mathfrak{f})^{\times}$. Let
$T=\{\beta\in S\colon(\mathrm{nr}(\beta)\mathcal{O}_{C}\bmod P_{\mathfrak{g}}^{+})\in\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})\}.$
Then $\Lambda$ has SFC if and only if $X_{\beta}$ is free over $\Lambda$ for every $\beta\in T$.
Proof.
By Proposition 8.3 (v),(vii) we have $T=\{\beta\in S\colon X_{\beta}\text{ is stably free over }\Lambda\}$. Thus the desired result follows easily from Theorem 8.7 and Corollary 8.5. ∎
8.2. An algorithm for determining SFC
The following algorithm determines whether an order has SFC or not. However, it is only practical for orders in algebras of ‘small’ dimension and in many situations, it is often much faster to use Algorithm 9.1 or 10.3.
Algorithm 8.9.
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$. If the following algorithm returns $\mathsf{true}$, then $\Lambda$ has SFC. If it returns $\mathsf{false}$, then $\Lambda$ fails SFC.
  1. (1)
    Compute the centre $C$ of $A$ and the decomposition of $A$ into simple $K$-algebras.
  2. (2)
    Compute a maximal $\mathcal{O}$-order $\mathcal{M}$ in $A$ containing $\Lambda$.
  3. (3)
    Check whether $\mathcal{M}$ has SFC. If not, return $\mathsf{false}$. If so and $\Lambda=\mathcal{M}$ return $\mathsf{true}$. Otherwise, continue with the steps below.
  4. (4)
    Compute $\mathfrak{f}:=\{x\in A\mid x\mathcal{M}\subseteq\Lambda\}\cap\{x\in A\mid\mathcal{M}x\subseteq\Lambda\}$.
  5. (5)
    Compute $\mathcal{O}_{C}:=\mathcal{M}\cap C$ and $\mathfrak{g}:=\mathfrak{f}\cap\mathcal{O}_{C}=\mathfrak{f}\cap C$.
  6. (6)
    Adjust $\mathfrak{f}$ (and thus $\mathfrak{g}$) to ensure that $\mathfrak{g}$ satisfies Assumption (G).
  7. (7)
    Compute a set of generators of $(\Lambda/\mathfrak{f})^{\times}$.
  8. (8)
    Compute $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$ and $\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$ as finite abelian groups.
  9. (9)
    Compute the set $(\mathcal{M}/\mathfrak{f})^{\times}$ and a set of representatives $S\subseteq\mathcal{M}\cap A^{\times}$ of $(\mathcal{M}/\mathfrak{f})^{\times}$.
  10. (10)
    Compute the set $T=\{\beta\in S\colon(\mathrm{nr}(\beta)\mathcal{O}_{C}\bmod P_{\mathfrak{g}}^{+})\in\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})\}$.
  11. (11)
    For each $\beta\in T$ check whether $X_{\beta}$ is free over $\Lambda$. If this is true for all $\beta\in T$, return $\mathsf{true}$. Otherwise return $\mathsf{false}$.
Proof of correctness.
In step (3), the correctness of an output $\mathsf{false}$ follows from Corollary 3.2 and the correctness of an output $\mathsf{true}$ is clear. If step (4) is reached then $\Lambda$ must be non-maximal. Note that $\mathfrak{f}$ as computed in step (4) is a two-sided ideal of $\mathcal{M}$ such that $\mathfrak{f}\mathcal{M}\subseteq\Lambda$, and so $\mathfrak{f}$ must be proper. In step (11), the correctness of the output follows from Theorem 8.7. ∎
Further details on each step.
Steps (1)–(5) can be performed using [FR85, 1.5 B], [Fri00, Kapitel 3 and 4], Corollary 3.6, [Fri00, Algorithmus (2.16)] and [Coh00, Algorithm 1.5.1], respectively.
Step (6): First note that for a given ideal $\mathfrak{g}$ there are only finitely many maximal ideals $\mathfrak{p}$ of $\mathcal{O}$ such that $\mathfrak{g}_{\mathfrak{p}}\neq\mathcal{O}_{C,\mathfrak{p}}$. If $\mathfrak{g}=\mathfrak{f}\cap\mathcal{O}_{C}$ fails Assumption (G) for $\mathfrak{P}\mid\mathfrak{p}$, then we replace $\mathfrak{f}$ by $\mathfrak{P}\mathfrak{f}$ and observe that $(\mathfrak{P}\mathfrak{f}\cap\mathcal{O}_{C})_{\mathfrak{P}}=\mathfrak{P}\mathfrak{f}_{\mathfrak{P}}\cap\mathcal{O}_{C,\mathfrak{P}}\subseteq\mathfrak{P}\left(\mathfrak{f}_{\mathfrak{P}}\cap\mathcal{O}_{C,\mathfrak{P}}\right)\subseteq\mathfrak{P}\mathcal{O}_{C,\mathfrak{P}}$.
Step (7) can be performed using the algorithm described in [BHJ22, §6.6] (this is a minor modification of [BB06, §3.4–3.7]).
Step (8): In the notation of §7.7, compute $\infty_{i}$ for $i=1,\ldots,r$ using [NS09, Theorem 3.5]. The structure of the finitely generated abelian group $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$ can then be computed component-wise using [Coh00, Algorithm 4.3.1]. Using the output of Step (7) and the definition of $\tilde{\nu}$, then determine generators of $\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$.
Steps (9) and (10): Since both rings $\mathcal{M}/\mathfrak{f}$ and $\Lambda/\mathfrak{f}$ are finite, their units can be computed as sets in finite time. (For a more practical approach, see §11). Given $\beta\in S$, algorithms for finitely generated abelian groups, as presented for example in [Coh00, §4.1], can be used to determine whether $\tilde{\nu}(\overline{\beta})\in\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$.
Step (11): Given $\beta\in T$, [BJ11, Algorithm 3.1] can be used to determine whether $X_{\beta}$ is free over $\Lambda$ (though stated for orders in group algebras, this algorithm can also be applied to orders in arbitrary finite-dimensional semisimple algebras as explained in [BJ11, Remark 4.9]). Note that this algorithm requires certain hypotheses, which ensure that one can find the preimage of the image of the canonical map ×$\mathcal{M}^{\times}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$, as well as a generator of $\mathcal{M}X_{\beta}$ as an $\mathcal{M}$-module. Since generators of $\mathcal{M}^{\times}$ can in principle be found using the algorithms of [BCNS15], and since $\mathcal{M}X_{\beta}=\mathcal{M}\beta$ by Proposition 8.3 (iii), these hypotheses are in fact unnecessary in this situation. ∎
Remark 8.10.
Steps (4)–(6) of Algorithm 8.9 can be replaced by the following steps.
  • (4)
    Compute $\mathcal{O}_{C}:=\mathcal{M}\cap C$ and $\mathfrak{g}:=\{x\in C\mid x\mathcal{M}\subseteq\Lambda\}$.
  • (5)
    If necessary, adjust $\mathfrak{g}$ to ensure that it satisfies Assumption (G).
  • (6)
    Compute $\mathfrak{f}:=\mathfrak{g}\mathcal{M}$.
Then by Lemma 7.11 the ideals $\mathfrak{f}$ and $\mathfrak{g}$ would satisfy $\mathfrak{g}=\mathfrak{f}\cap\mathcal{O}_{C}=\mathfrak{f}\cap C$. However, the disadvantage of this approach is that in general it will produce a smaller choice of $\mathfrak{f}$ and thus a larger choices of the sets $S$ and $T$.
Remark 8.11.
Algorithm 8.9 can be modified to use Corollary 8.8 instead of Theorem 8.7. This involves replacing step (9) by
  • (9)
    Compute a set of representatives $S\subseteq\mathcal{M}\cap A^{\times}$ of ×\$\mathcal{M}^{\times}\backslash(\mathcal{M}/\mathfrak{f})^{\times}{}{/}{}(\Lambda/\mathfrak{f})^{\times}$.
In principle, this version of the algorithm should be much more efficient since the sets $S$ and $T$ become much smaller. However, Step (9) itself comes with its own challenges.
9. Orders without SFC
9.1. An algorithm to show that an order fails SFC
The following algorithm can be viewed as a modified version of Algorithm 8.9 that cannot prove that an order has SFC, but can be used to show that it fails SFC.
Algorithm 9.1.
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$. If the following algorithm returns $\mathsf{false}$, then $\Lambda$ fails SFC. If it returns $\mathsf{true}$, then $\Lambda$ is maximal and has SFC. If it returns $\mathsf{inconclusive}$, then nothing can be said about whether $\Lambda$ has SFC or not.
  1. (1)
    Perform steps (1) and (2) of Algorithm 8.9.
  2. (2)
    Check whether $\mathcal{M}$ has SFC. If not, return $\mathsf{false}$. If so and $\Lambda=\mathcal{M}$ return $\mathsf{true}$. Otherwise, continue with the steps below.
  3. (3)
    Perform steps (4)–(8) of Algorithm 8.9.
  4. (4)
    Choose $r\geq 1$ and compute ‘random’ elements $\beta_{1},\dotsc,\beta_{r}\in\mathcal{M}\cap A^{\times}$ such that $\overline{\beta_{i}}\in(\mathcal{M}/\mathfrak{f})^{\times}$ and $(\mathrm{nr}(\beta_{i})\mathcal{O}_{C}\bmod P_{\mathfrak{g}}^{+})\in\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$ for all $i\in\{1,\ldots,r\}$.
  5. (5)
    For each $i\in\{1,\ldots,r\}$ check whether $X_{\beta_{i}}$ is free over $\Lambda$. If this is true for all $i\in\{1,\ldots,r\}$, return $\mathsf{inconclusive}$. Otherwise, return $\mathsf{false}$.
Proof of correctness.
In step (2), the correctness of an output $\mathsf{false}$ follows from Corollary 3.2 and the correctness of an output $\mathsf{true}$ is clear. Step (4) ensures that $X_{\beta_{i}}$ is stably free over $\Lambda$ for each $i\in\{1,\ldots,r\}$ by Proposition 8.3 (v),(vii). Thus output of step (5) is correct because if there exists $i\in\{1,\ldots,r\}$ such that $X_{\beta_{i}}$ is not free over $\Lambda$, then $\Lambda$ clearly fails SFC. Note that step (2) is in fact optional, in that the algorithm is still correct even if it is omitted. ∎
Further details on each step.
Steps (1)–(3) and step (5) of Algorithm 9.1 can be performed in the same way as the corresponding steps of Algorithm 8.9. To find the ‘random’ elements in step (4), one can determine elements of $\mathcal{M}\cap A^{\times}$ with random coordinates in a fixed basis of bounded height and then check whether they are invertible modulo $\mathfrak{f}$. Alternatively, one could also determine elements of $\mathcal{M}\cap A^{\times}$ generating $(\mathcal{M}/\mathfrak{f})^{\times}$ using the algorithm described in [BHJ22, §6.6] (which is a minor modification of [BB06, §3.4–3.7]) and then take random words of bounded length in the generators. ∎
Remark 9.2.
Thanks to the fact that Algorithm 9.1 does not need to perform steps (9) or (10) of Algorithm 8.9, in practice the former is often much faster than the latter if one wants to prove that a given order $\Lambda$ fails SFC.
Remark 9.3.
Due to the invocation of the (expensive) freeness test of [BJ11, Algorithm 3.1], Algorithm 9.1 is only practical when $\Lambda$ is an order in an algebra $A$ of ‘moderate’ dimension. However, it is also possible to obtain results when $A$ is of larger dimension as follows. A projection of $K$-algebras $A\to B$ induces a projection of $\Lambda$ onto an $\mathcal{O}$-order $\Gamma$ in $B$. If we can apply Algorithm 9.1 to show that $\Gamma$ fails SFC, then Theorem 3.1 shows that $\Lambda$ also fails SFC. Since we are free to choose $B$, we can take $B$ to be a direct sum of simple components of $A$ of small dimension, chosen such that all non-Eichler simple components of $A$ are also components of $B$ and Algorithm 9.1 finishes quickly.
9.2. Integral group rings without SFC
We now use Algorithm 9.1 and Remark 9.3 to show that certain integral group rings fail SFC. Parts of the following result have already proven by Swan [Swa83] and Chen [Che86] (see Remark 9.5).
Theorem 9.4.
Let $G$ be one of the following finite groups:
  1. (i)
    $Q_{4n}\times C_{2}$ for $2\leq n\leq 5$,
  2. (ii)
    $\tilde{T}\times C_{2}^{2}$, $\tilde{O}\times C_{2}$, $\tilde{I}\times C_{2}$, or
  3. (iii)
    $G_{(32,14)}$, $G_{(36,7)}$, $G_{(64,14)}$, $G_{(100,7)}$.
Let $e$ be the central idempotent of $\mathbb{Q}[G]$ such that $e\mathbb{Q}[G]$ is the direct sum of all simple components of $\mathbb{Q}[G]$ that are either commutative or totally definite quaternion algebras. Then $e\mathbb{Z}[G]$ fails SFC. Moreover, $\mathbb{Z}[G]$ fails SFC but $\mathbb{Z}[G/N]$ has SFC for every proper quotient $G/N$.
Remark 9.5.
Theorem 9.4 agrees with previously known results. Recall from §4 that Swan [Swa83, Theorem 15.1] already proved that $\mathbb{Z}[Q_{8}\times C_{2}]$ fails SFC and Chen [Che86] already proved that $\mathbb{Z}[G]$ fails SFC for $G=Q_{12}\times C_{2}$, $Q_{16}\times C_{2}$, $Q_{20}\times C_{2}$, $G_{(36,7)}$, $G_{(100,7)}$.
Proof of Theorem 9.4.
For each group $G$ listed, verification that $e\mathbb{Z}[G]$ fails SFC was performed using Algorithm 9.1. Theorem 3.1 then implies that $\mathbb{Z}[G]$ also fails SFC.
We now show that $\mathbb{Z}[G/N]$ has SFC for every proper quotient $G/N$. By Lemma 4.1, it suffices to consider maximal proper quotients $H=G/N$. If $\mathbb{Q}[H]$ satisfies the Eichler condition, then $\mathbb{Z}[H]$ has LFC by Theorem 3.4, and if $H$ is isomorphic to any of
$Q_{8},\,Q_{12},\,Q_{16},\,Q_{20},\,\tilde{T},\,\tilde{O},\,\tilde{I},$ (9.1)
then $\mathbb{Z}[H]$ has LFC by Theorem 4.2. Therefore it suffices to consider maximal quotients $H$ such that $\mathbb{Q}[H]$ does not satisfy the Eichler condition and is not isomorphic to any of the groups in (9.1). The only groups $G$ listed that have such maximal quotients are $\tilde{T}\times C_{2}^{2}$, $G_{(32,14)}$ and $G_{(64,14)}$. For $G=\tilde{T}\times C_{2}^{2}$, each such $H$ is isomorphic to $\tilde{T}\times C_{2}$, and so $\mathbb{Z}[H]$ has SFC by Theorem 6.6. For $G=G_{(32,14)}$, each such $H$ is isomorphic to $G_{(16,4)}\cong C_{4}\rtimes C_{4}$ and so $\mathbb{Z}[H]$ has SFC by Theorem 4.11. For $G=G_{(64,14)}$, each such $H$ is isomorphic to $G_{(32,10)}\cong Q_{8}\rtimes C_{4}$. Since $G_{(32,10)}$ has a quotient isomorphic to $Q_{16}$ and $m_{\mathbb{H}}(G_{(32,10)})=m_{\mathbb{H}}(Q_{16})=2$, the group ring $\mathbb{Z}[G_{(32,10)}]$ has SFC by Theorem 4.10. ∎
Remark 9.6.
Let $G$ be a group listed in Theorem 9.4. Then Corollary 4.4 implies that every maximal $\mathbb{Z}$-order $\mathcal{M}$ in $\mathbb{Q}[G]$ has LFC. Thus if such an $\mathcal{M}$ contains $\mathbb{Z}[G]$, then $e\mathcal{M}$ is a maximal $\mathbb{Z}$-order containing $e\mathbb{Z}[G]$, and by Theorem 3.1 $e\mathcal{M}$ also has LFC (and thus SFC). Moreover, $e\mathbb{Z}[G]$ is not maximal. Therefore it is not necessary to perform Step (2) of Algorithm 9.1 when applied to $e\mathbb{Z}[G]$.
The following is the most general result to date on integral group rings that fail SFC.
Corollary 9.7.
Let $G$ be a finite group with a quotient isomorphic to a group listed in Theorem 9.4 or to $Q_{4n}$ for some $n\geq 6$. Then $\mathbb{Z}[G]$ fails SFC.
Proof.
This follows from Lemma 4.1, Theorem 4.2 and Theorem 9.4. ∎
Remark 9.8.
The following question naturally arises: for a finite group $G$ and $e$ as in the statement of Theorem 9.4, is it always true that $\mathbb{Z}[G]$ has SFC if and only if $e\mathbb{Z}[G]$ has SFC? One direction is clear by Theorem 3.1. If the other direction could be proven, then this would make the problem of determining whether $\mathbb{Z}[G]$ has SFC much easier from an algorithmic point of view. Of course, one can also ask the analogous question for LFC.
Remark 9.9.
Let $G$ be a finite group and let $\tilde{e}$ be the central idempotent of $\mathbb{Q}[G]$ such that $\tilde{e}\mathbb{Q}[G]$ is the direct sum of all simple components of $\mathbb{Q}[G]$ that are totally definite quaternion algebras. Then for $G=Q_{16}\times C_{2}$, the order $\tilde{e}\mathbb{Z}[G]$ fails SFC, but for $G=Q_{12}\times C_{2}$, it does have SFC. Thus $e$ cannot be replaced by $\tilde{e}$ for all groups listed in Theorem 9.4.
10. Determining SFC using fibre products
The main result of this section is Algorithm 10.3, which may be viewed as a constructive version of Corollary 5.12.
10.1. Setup
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$. Let $I$ and $J$ be two-sided ideals of $\Lambda$ such that $I\cap J=0$ and $K(I+J)=A$. Let $\Lambda_{1}=\Lambda/I$, let $\Lambda_{2}=\Lambda/J$ and let $\overline{\Lambda}=\Lambda/(I+J)$. Then $\Lambda_{1}$ and $\Lambda_{2}$ are $\mathcal{O}$-orders and $\overline{\Lambda}$ is an $\mathcal{O}$-torsion $\mathcal{O}$-algebra. Then as in Example 5.1, we have a fibre product
${\Lambda}$${\Lambda_{1}}$${\Lambda_{2}}$${\overline{\Lambda},}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{g_{1}}$$\scriptstyle{g_{2}}$ (10.1)
where each of $\pi_{1},\pi_{2},g_{1},g_{2}$ is the relevant canonical surjection.
Let $A_{1}:=K\Lambda_{1}\cong KJ$ and let $A_{2}:=K\Lambda_{1}\cong KI$. Then by making the appropriate identifications, we can and do assume that
$\Lambda=\{(x_{1},x_{2})\in\Lambda_{1}\oplus\Lambda_{2}\mid g_{1}(x_{1})=g_{2}(x_{2})\}\subseteq\Lambda_{1}\oplus\Lambda_{2}\subseteq A_{1}\oplus A_{2}=A.$ (10.2)
In particular, $\pi_{i}$ is the map induced by the canonical projection $A\to A_{i}$ for $i=1,2$.
Recall from Proposition 5.2 that for $u\in\overline{\Lambda}^{\times}$, we define
$M(u):=\{(x_{1},x_{2})\in\Lambda_{1}\oplus\Lambda_{2}\mid g_{1}(x_{1})u=g_{2}(x_{2})\}.$
In particular, this is a locally free $\Lambda$-lattice of rank $1$. We refer to such a lattice $M(u)$ as a Reiner–Ullom module. Recall from Corollary 5.3 that we have a group homomorphism
$\partial:\overline{\Lambda}^{\times}\longrightarrow\mathrm{Cl}(\Lambda),\quad u\mapsto[M(u)]_{\mathrm{st}}.$
We shall relate Reiner–Ullom modules to generalised Swan modules and use this to give an alternative description of $\ker(\partial)$. Under the further assumption that $A_{2}$ is Eichler, this will lead to an algorithm to test whether $\Lambda$ has SFC or not.
10.2. Generalised Swan modules and Reiner–Ullom modules
For $i=1,2$, let $\mathcal{M}_{i}$ be a choice of $\mathcal{O}$-order in $A_{i}$ containing $\Lambda_{i}$ (such choices exists by [Rei03, (10.4)]) and let $\mathfrak{f}_{i}$ be any proper full two-sided ideal of $\mathcal{M}_{i}$ that is contained in $\Lambda\cap A_{i}$. Let $\mathcal{M}=\mathcal{M}_{1}\oplus\mathcal{M}_{2}$ and let $\mathfrak{f}=\mathfrak{f}_{1}\oplus\mathfrak{f}_{2}$. For $i=1,2$, let $C_{i}$ be the centre of $A_{i}$ and let $\mathfrak{g}_{i}:=\mathfrak{f}_{i}\cap\mathcal{O}_{C_{i}}=\mathfrak{f}_{i}\cap C_{i}$. Let $C=C_{1}\oplus C_{2}$ and let $\mathfrak{g}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{2}$. We set
$\mathfrak{h}_{1}:=\ker(g_{1}),\quad\mathfrak{h}_{2}:=\ker(g_{2}),\quad\mathfrak{h}:=\pi_{1}^{-1}(\mathfrak{h}_{1})=\pi_{2}^{-1}(\mathfrak{h}_{2})=I+J,$
and observe that
$\overline{\Lambda}=\Lambda/\mathfrak{h}\cong\Lambda_{1}/\mathfrak{h}_{1}\cong\Lambda_{2}/\mathfrak{h}_{2}.$
Using the identification in (10.2) we have $\mathfrak{h}_{i}=\Lambda\cap A_{i}$ for $i=1,2$.
Since for any $x_{1}\in\Lambda_{1}$ with $(x_{1},0)\in\Lambda$ we have $x_{1}\in\ker(g_{1})$, it follows that $\mathfrak{f}_{1}\subseteq\mathfrak{h}_{1}$. Similarly, $\mathfrak{f}_{2}\subseteq\mathfrak{h}_{2}$ and therefore $\mathfrak{f}\subseteq\mathfrak{h}$. Recall from Definition 8.1 that for $\beta\in\mathcal{M}\cap A^{\times}$, the generalised Swan module $X_{\beta}=X_{\beta,\Lambda,\mathcal{M}}$ is defined to be $\mathcal{M}\beta\cap\Lambda$.
The following result shows that under mild assumptions, Reiner–Ullom modules are in fact (isomorphic to) generalised Swan modules.
Proposition 10.1.
Suppose that $\mathfrak{g}$ satisfies Assumption (G). Let $\beta=(\beta_{1},\beta_{2})\in\Lambda_{1}\oplus\Lambda_{2}$ with $\beta\in A^{\times}$ and $\overline{\beta_{i}}\in\left(\Lambda_{i}/\mathfrak{f}_{i}\right)^{\times}$ for $i=1,2$. Let $u=g_{1}(\beta_{1})g_{2}(\beta_{2})^{-1}\in\overline{\Lambda}^{\times}$. Then $X_{\beta}=M(u)\beta$, so in particular, $X_{\beta}\cong M(u)$ as $\Lambda$-lattices.
Proof.
Let $\mathfrak{p}$ be a maximal ideal of $\mathcal{O}$. By [Rei03, Theorem 4.21], it suffices to prove that $(X_{\beta})_{\mathfrak{p}}=(M(u)\beta)_{\mathfrak{p}}$. Note that $(M(u)\beta)_{\mathfrak{p}}=M(u)_{\mathfrak{p}}\beta$.
Suppose that $\Lambda_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}$. Since $\Lambda_{\mathfrak{p}}\subseteq\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}$ and $\Lambda_{\mathfrak{p}}$ is maximal, we have that $\Lambda_{\mathfrak{p}}=\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}$. Thus localising the fibre product (5.1) at $\mathfrak{p}$ implies that $\mathfrak{h}_{i,\mathfrak{p}}=\mathcal{M}_{i,\mathfrak{p}}=\Lambda_{i,\mathfrak{p}}$ for $i=1,2$ and $\overline{\Lambda}_{\mathfrak{p}}=0$. It is then easy to see that $M(u)_{\mathfrak{p}}=\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}$ and hence $M(u)_{\mathfrak{p}}\beta=\mathcal{M}_{\mathfrak{p}}\beta=(X_{\beta})_{\mathfrak{p}}$, where the last equality follows from Lemma 8.2.
Now suppose that $\Lambda_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$. Then $\beta\in\mathcal{M}_{\mathfrak{p}}^{\times}$ by Lemma 8.2. Thus $\beta_{i}\in\mathcal{M}_{i,\mathfrak{p}}^{\times}$ and so in fact $\beta_{i}\in\Lambda_{i,\mathfrak{p}}\cap\mathcal{M}_{i,\mathfrak{p}}^{\times}=\Lambda_{i,\mathfrak{p}}^{\times}$ (see [Kar92, Chapter I, Theorem 7.9(iii)]). Then we have
$M(u)_{\mathfrak{p}}\beta$ =$=\{(a_{1}\beta_{1},a_{2}\beta_{2})\mid(a_{1},a_{2})\in\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}\text{ and }g_{1,\mathfrak{p}}(a_{1})u=g_{2,\mathfrak{p}}(a_{2})\}$
=$=\{(a_{1}\beta_{1},a_{2}\beta_{2})\mid(a_{1},a_{2})\in\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}\text{ and }g_{1,\mathfrak{p}}(a_{1}\beta_{1})=g_{2,\mathfrak{p}}(a_{2}\beta_{2})\}$
=$=\{(b_{1},b_{2})\mid(b_{1},b_{2})\in\Lambda_{1,\mathfrak{p}}\oplus\Lambda_{2,\mathfrak{p}}\text{ and }g_{1,\mathfrak{p}}(b_{1})=g_{2,\mathfrak{p}}(b_{2})\}$
$=\Lambda_{\mathfrak{p}}=(X_{\beta})_{\mathfrak{p}},$
where the last equality follows from Lemma 8.2. ∎
10.3. Alternative description of $\ker(\partial)$
Recall from (7.8) and (7.9) respectively that we have an exact sequence
$\left(\Lambda/\mathfrak{f}\right)^{\times}\stackrel{{\scriptstyle\nu}}{{\longrightarrow}}I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}\stackrel{{\scriptstyle\mu}}{{\longrightarrow}}\mathrm{Cl}(\Lambda)\longrightarrow 0,$
and that the reduced norm map induces a group homomorphism
$\tilde{\nu}:(\mathcal{M}/\mathfrak{f})^{\times}\longrightarrow I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$
that extends $\nu$. Therefore $(I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+})/\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})\cong\mathrm{Cl}(\Lambda)$. Let
$\tilde{\mu}:I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}\longrightarrow(I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+})/\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$
denote the canonical surjection. Define
$\nu_{2}^{\prime}\colon(\Lambda_{2}/\mathfrak{h}_{2})^{\times}\longrightarrow(I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+})/\tilde{\nu}((\Lambda/\mathfrak{f})^{\times}),\quad x_{2}\mapsto\tilde{\mu}(\tilde{\nu}((\overline{1},\overline{\alpha_{2}}))),$ (10.3)
where $\alpha_{2}\in\Lambda_{2}\cap A_{2}^{\times}$ is any element such that $\overline{\alpha_{2}}\in(\Lambda_{2}/\mathfrak{f}_{2})^{\times}$ and $x_{2}=(\alpha_{2}\bmod\mathfrak{h}_{2})$.
Lemma 10.2.
The map $\nu_{2}^{\prime}$ is a well-defined group homomorphism. If $\mathfrak{g}$ satisfies Assumption (G) then $\ker(\partial)={\tilde{g}_{2}}(\ker(\nu^{\prime}_{2}))$, where $\tilde{g}_{2}:\Lambda_{2}/\mathfrak{h}_{2}\cong\overline{\Lambda}$ is the ring isomorphism induced by $g_{2}$.
Proof.
Let $\iota_{2}\colon(\Lambda_{2}/\mathfrak{f}_{2})^{\times}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$ and $\theta_{2}\colon(\Lambda_{2}/\mathfrak{f}_{2})^{\times}\longrightarrow(\Lambda_{2}/\mathfrak{h}_{2})^{\times}$ be the canonical group homomorphisms (recall from §10.2 that $\mathfrak{f}_{2}\subseteq\mathfrak{h}_{2}$). Then $\iota_{2}$ is clearly injective and $\theta_{2}$ is surjective by [Bas68, Chapter III, (2.9) Corollary]. Moreover, it follows from the definition of $\nu_{2}^{\prime}$ that $\tilde{\mu}\circ\tilde{\nu}\circ\iota_{2}=\theta_{2}\circ\nu_{2}^{\prime}$. Every element in $\ker(\theta_{2})$ is represented by an element of $\Lambda_{2}$ of the form $\lambda_{2}=1+h_{2}$ with $h_{2}\in\mathfrak{h}_{2}$. Since
$(1,\lambda_{2})=(1,1)+(0,h_{2})=1+(0,h_{2})\in\Lambda,$
we have $\tilde{\nu}(\overline{1},\overline{\lambda_{2}})\in\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$. Therefore $\nu_{2}^{\prime}$ is a well-defined group homomorphism.
Let $x_{2}\in\left(\Lambda_{2}/\mathfrak{h}_{2}\right)^{\times}$ and let $\alpha_{2}$ be as in the statement of the lemma. Then we have
$x_{2}\in\ker(\nu_{2}^{\prime})$ $\Longleftrightarrow\ \tilde{\nu}((\overline{1},\overline{\alpha_{2}}))\in\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$
$\Longleftrightarrow\ [X_{(1,\alpha_{2})}]_{\mathrm{st}}=0$ (Proposition 8.3 (vii))
$\Longleftrightarrow\ [M(g_{2}(\alpha_{2})^{-1})]_{\mathrm{st}}=0$ (Proposition 10.1)
$\Longleftrightarrow\ \tilde{g}_{2}(\alpha_{2}\bmod\mathfrak{h}_{2})\in\ker(\partial).$
Since $x_{2}=(\alpha_{2}\bmod\mathfrak{h}_{2})$ in $\Lambda_{2}/\mathfrak{h}_{2}\cong\overline{\Lambda}$, we conclude that $\ker(\partial)=\tilde{g}_{2}(\ker({\nu}_{2}^{\prime}))$. ∎
10.4. An algorithm for determining SFC using fibre products
The following algorithm is based on Corollary 5.12.
Algorithm 10.3.
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$ with a fibre product as in (10.1). Assume that $A_{1}$ is nontrivial and satisfies the Eichler condition. If the following algorithm returns $\mathsf{true}$, then $\Lambda$ has SFC. If it returns $\mathsf{false}$, then $\Lambda$ fails SFC.
  1. (1)
    Compute the centre $C=C_{1}\oplus C_{2}$ of $A=A_{1}\oplus A_{2}$ and the decomposition of $A=A_{1}\oplus A_{2}$ into simple $K$-algebras.
  2. (2)
    For $i=1,2$, compute a maximal $\mathcal{O}$-order $\mathcal{M}_{i}$ in $A_{i}$ containing $\Lambda_{i}$.
  3. (3)
    Check whether $\mathcal{M}_{2}$ has SFC. If not, return $\mathsf{false}$. If so, continue with the steps below.
  4. (4)
    For $i=1,2$, compute $\mathfrak{f}_{i}:=\{x\in A_{i}\mid x\mathcal{M}_{i}\subseteq\Lambda\}\cap\{x\in A_{i}\mid\mathcal{M}_{i}x\subseteq\Lambda_{i}\}$.
  5. (5)
    For $i=1,2$, compute $\mathcal{O}_{C_{i}}:=\mathcal{M}_{2}\cap C_{i}$ and $\mathfrak{g}_{i}:=\mathfrak{f}_{i}\cap\mathcal{O}_{C_{i}}=\mathfrak{f}_{i}\cap C_{i}$.
  6. (6)
    For $i=1,2$, adjust $\mathfrak{f}_{i}$ (and thus $\mathfrak{g}_{i}$) to ensure that $\mathfrak{g}_{i}$ satisfies Assumption (G).
  7. (7)
    Check whether $\Lambda_{2}$ has SFC. If not, return $\mathsf{false}$. If so, continue with the steps below.
  8. (8)
    Compute $\mathfrak{h}_{2}=\ker(g_{2})$ and $\overline{\Lambda}_{2}=\Lambda_{2}/\mathfrak{h}_{2}$.
  9. (9)
    Compute $\overline{\Lambda}_{2}^{\times}$ as a finite group.
  10. (10)
    Compute $\overline{\Lambda}_{2}^{\times\mathrm{ab}}=\overline{\Lambda}_{2}^{\times}{}{/}{}[\overline{\Lambda}_{2}^{\times},\overline{\Lambda}_{2}^{\times}]$ as a finite abelian group.
  11. (11)
    Compute a set of generators of $(\Lambda/\mathfrak{f})^{\times}$, where $\mathfrak{f}=\mathfrak{f}_{1}\oplus\mathfrak{f}_{2}$.
  12. (12)
    Compute $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$ and $\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$ as finite abelian groups, where $\mathfrak{g}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{2}$.
  13. (13)
    Compute generators of the subgroup $\ker(\nu_{2}^{\prime})\leq\overline{\Lambda}_{2}^{\times}$.
  14. (14)
    Compute elements $\overline{\gamma}_{1},\ldots,\overline{\gamma}_{r}\in\overline{\Lambda}_{2}^{\times}$ such that their images are generators of the subgroup $\ker(\nu_{2}^{\prime})/[\overline{\Lambda}_{2}^{\times},\overline{\Lambda}_{2}^{\times}]\leq\overline{\Lambda}_{2}^{\times\mathrm{ab}}$.
  15. (15)
    For each $i\in\{1,\ldots,r\}$, choose an element $\beta_{i}\in\Lambda_{2}\cap A_{2}^{\times}$ with $(\beta_{i}\bmod\mathfrak{f}_{2})\in\left(\Lambda_{2}/\mathfrak{f}_{2}\right)^{\times}$ and such that $\overline{\gamma}_{i}=(\beta_{i}\bmod{\mathfrak{h}_{2}})$.
  16. (16)
    For each $i\in\{1,\ldots,r\}$, check whether $X_{(1,\beta_{i})}$ is free over $\Lambda$. If this is true for all $i\in\{1,\ldots,r\}$, return $\mathsf{true}$. Otherwise, return $\mathsf{false}$.
Proof of correctness.
By Theorem 3.4 and the assumption that $A_{1}$ satisfies the Eichler condition, $\Lambda_{1}$ must have SFC. Thus by Corollary 3.3, $\Lambda_{1}\oplus\Lambda_{2}$ has SFC if and only if $\Lambda_{2}$ has SFC. Therefore the correctness of an output $\mathsf{false}$ in either step (3) or step (7) follows from Corollary 3.2 since $\Lambda\subseteq\Lambda_{1}\oplus\Lambda_{2}\subseteq\Lambda_{1}\oplus\mathcal{M}_{2}$.
By Lemma 10.2, we have $\ker(\partial)={\tilde{g}_{2}}(\ker(\nu^{\prime}_{2}))$, where $\tilde{g}_{2}:\Lambda_{2}/\mathfrak{h}_{2}\cong\overline{\Lambda}$ is the ring isomorphism induced by $g_{2}$. Thus $\hat{g}_{2}^{-1}(\overline{\gamma}_{1}),\ldots,\hat{g}_{2}^{-1}(\overline{\gamma}_{r})$ is a set of generators of $\ker(\partial)/[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]$. Moreover, by Proposition 10.1, for each $i\in\{1,\ldots,r\}$, we have $X_{(1,\beta_{i})}\cong M(g_{2}(\beta_{i})^{-1})$ as $\Lambda$-lattices. Therefore the correctness of the output in step (16) follows from Corollary 5.12 with $N=[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]$. ∎
Further details on each step.
Steps (1)–(6), (11), and (12) can be performed in the same way as the corresponding steps of Algorithm 8.9.
In principle, Step (7) can be performed using any appropriate algorithm, such as Algorithm 8.9, for example. Note that Algorithm 10.3 itself can be applied recursively, which is useful because the choice of fibre product can have a large impact on the running time.
Step (8) is straightforward and Step (9) can be performed using the methods outlined in §11.3. Note that, in practice, Step (9) is often the bottleneck in the computation. Steps (10), (13), and (14) can be performed using standard methods in computational group theory, as described in [HEO05], for example, and the explicit definition of $\nu_{2}^{\prime}$ given in (10.3). Step (15) is straightforward and in Step (16), the freeness testing can be performed either as described in Step (11) of Algorithm 8.9 or the (more efficient) method described in §12. ∎
11. Primary decomposition and unit groups of quotient rings
11.1. Setup
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$ and let $\mathfrak{h}$ be any proper full two-sided ideal of $\Lambda$. In §11.2, we describe a practical algorithm to compute a set of generators of the unit group $(\Lambda/\mathfrak{h})^{\times}$. In §11.3, we then describe a practical algorithm to compute $(\Lambda/\mathfrak{h})^{\times}$ as an abstract group. These are important subtasks in Algorithm 10.3.
11.2. Primary decomposition and generators of unit groups of quotient rings
Let $C$ be an étale $K$-algebra and let $\Gamma$ be an $\mathcal{O}$-order in $C$. Let $\mathfrak{g}$ be a proper ideal of $\Gamma$ of full rank. We now describe the computation of a primary decomposition of $\mathfrak{g}$, that is, the computation of coprime primary ideals $\mathfrak{q}_{1},\dotsc,\mathfrak{q}_{r}$ of $\Gamma$, such that
$\mathfrak{g}=\prod_{i=1}^{r}\mathfrak{q}_{i}=\bigcap_{i=1}^{r}\mathfrak{q}_{i}.$
Proposition 11.1.
Let $C$ be an étale $K$-algebra and let $\Gamma$ be an $\mathcal{O}$-order in $C$. Let $\mathfrak{g}$ be a proper ideal of $\Gamma$ of full rank. Let $\mathfrak{p}_{1},\dotsc,\mathfrak{p}_{r}$ be the maximal ideals of $\Gamma$ containing $\mathfrak{g}$. Then the following hold.
  1. (i)
    There exist exponents $e_{1},\ldots,e_{r}\in\mathbb{Z}_{>0}$ such that
    $\mathfrak{p}_{1}^{e_{1}}\cdots\mathfrak{p}_{r}^{e_{r}}\subseteq\mathfrak{g}\subseteq\mathfrak{p}_{1}\cdots\mathfrak{p}_{r}.$
  2. (ii)
    For each $i\in\{1,\dotsc,r\}$, the ideal $\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g}$ is $\mathfrak{p}_{i}$-primary.
  3. (iii)
    A primary decomposition of $\mathfrak{g}$ is given by
    $\mathfrak{g}=\bigcap_{i=1}^{r}(\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g})=\prod_{i=1}^{r}(\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g}).$
Proof.
(i) As in the proof of [Neu99, (3.4) Lemma] one can show that there is a product of maximal ideals $\mathfrak{q}_{i}$ of $\Gamma$ such that $\mathfrak{q}_{1}\cdots\mathfrak{q}_{m}\subseteq\mathfrak{g}$. Since $\mathfrak{q}_{1}\cdots\mathfrak{q}_{m}\subseteq\mathfrak{g}\subseteq\mathfrak{p}_{i}$ for $i=1,\ldots,r$ each of the maximal ideals $\mathfrak{q}_{i}$ is one of $\mathfrak{p}_{1},\ldots,\mathfrak{p}_{r}$. Hence there exist $e_{1},\ldots,e_{r}\in\mathbb{Z}_{>0}$ such that $\mathfrak{p}_{1}^{e_{1}}\cdots\mathfrak{p}_{r}^{e_{r}}\subseteq\mathfrak{g}$.
(ii) Since $\mathfrak{p}_{i}^{e_{i}}\subseteq\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g}\subseteq\mathfrak{p}_{i}$, the radical of $\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g}$ is the maximal ideal $\mathfrak{p}_{i}$. This implies that $\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g}$ is $\mathfrak{p}_{i}$-primary ([AM69, Prop. (4.2)]).
(iii) It suffices to show that $\mathfrak{g}=\bigcap_{i=1}^{r}(\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g})=\prod_{i=1}^{r}(\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g})$. The second equality is clear because the ideals $\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g}$ are pairwise coprime. For the first equality the inclusion $\subseteq$” is obvious. Conversely, $\prod_{i=1}^{r}(\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g})\subseteq\mathfrak{p}_{1}^{e_{1}}\cdots\mathfrak{p}_{r}^{e_{r}}+\mathfrak{g}=\mathfrak{g}$. ∎
Remark 11.2.
Let $\mathcal{O}_{C}$ be the (unique) maximal $\mathcal{O}$-order in $C$. This is equal to the integral closure of $\Gamma$ in $C$. By the going-up theorem each of the maximal ideals $\mathfrak{p}_{i}$ is of the form $\mathfrak{P}_{i}\cap\Gamma$ for a maximal ideal $\mathfrak{P}_{i}$ of $\mathcal{O}_{C}$. Let $\{\mathfrak{P}_{1},\ldots,\mathfrak{P}_{n}\}$ be the set of maximal ideals of $\mathcal{O}_{C}$ with $\mathfrak{g}\mathcal{O}_{C}\subseteq\mathfrak{P}_{i}$. Then
$\{\mathfrak{p}_{1},\ldots,\mathfrak{p}_{r}\}=\{\mathfrak{P}_{1}\cap\Gamma,\ldots,\mathfrak{P}_{n}\cap\Gamma\}.$
Note that $\mathfrak{P}_{1},\ldots,\mathfrak{P}_{n}$ can be computed using standard methods of algorithmic number theory [Len92, Theorem 4.9].
Corollary 11.3.
Let $C$ be be an étale $K$-algebra and let $\Gamma$ be an $\mathcal{O}$-order in $C$. Let $\mathfrak{g}$ be non-trivial proper ideal of $\Gamma$ of full rank. Then there exists an algorithm that computes the primary decomposition of $\mathfrak{g}$.
Corollary 11.4.
Let $\Lambda$ be an $\mathcal{O}$-order in a finite-dimensional semisimple $K$-algebra $A$ and let $\mathfrak{h}$ be a proper full two-sided ideal of $\Lambda$. Then there exists an algorithm that determines generators of $(\Lambda/\mathfrak{h})^{\times}$.
Proof.
Let $C$ denote the centre of $A$ and let $\mathcal{O}_{C}$ denote the (unique) maximal $\mathcal{O}$-order contained in $C$. Let $\mathfrak{g}=\mathfrak{h}\cap C$. Then $\mathfrak{g}$ is an ideal of an $\mathcal{O}$-order $\Gamma$, which may be strictly contained in $\mathcal{O}_{C}$. In the case that $\mathfrak{h}$ is equal to $\mathfrak{f}$, a two-sided ideal of a maximal $\mathcal{O}$-order $\mathcal{M}$ in $A$ containing $\Lambda$, the desired algorithm is described in [BHJ22, §6.6] (this is a minor modification of [BB06, §3.4–3.7]). This relies crucially on the computation of the primary decomposition of $\mathfrak{g}$, which can be accomplished using work of Bley–Endres [BE05] under the assumption that $\Gamma=\mathcal{O}_{C}$ (which must hold when $\mathfrak{h}=\mathfrak{f}$). In fact, the computation of the primary decomposition of $\mathfrak{g}$ is the only obstacle to generalising the algorithm to arbitrary choices of $\mathfrak{h}$. By removing the restriction that $\Gamma=\mathcal{O}_{C}$ on the computation of primary decompositions in Corollary 11.3, we thereby obtain an algorithm for the computation of generators of $(\Lambda/\mathfrak{h})^{\times}$ for arbitrary proper full two-sided ideals $\mathfrak{h}$ of $\Lambda$. ∎
11.3. Group structure of unit groups of finite rings
Let $\overline{\Lambda}=\Lambda/\mathfrak{h}$. Then $\overline{\Lambda}$ is a finite ring and an important challenge to compute its units $\overline{\Lambda}^{\times}$ as an abstract group. The point is that, although we can compute generators of $\overline{\Lambda}^{\times}$ by Corollary 11.4, we require a representation of $\overline{\Lambda}^{\times}$ that will allow us to efficiently determine the group-theoretic properties required for applications, such as the computation of the abelianisation $\overline{\Lambda}^{\times\mathrm{ab}}$. Specifically, representations which are useful for actual computations include realisations of $\overline{\Lambda}^{\times}$ as a subgroup of either $\mathrm{GL}_{n}(\mathbb{F}_{q})$ for some finite field $\mathbb{F}_{q}$, or of symmetric groups $\mathfrak{S}_{d}$, both of course with reasonable sizes for $n$ and $\lvert k\rvert$, or $d$, respectively (see [HEO05]).
Let $\mathrm{Aut}_{\mathbb{Z}}(\overline{\Lambda})$ denote the group of $\mathbb{Z}$-linear automorphisms of the (additive) $\mathbb{Z}$-module $\overline{\Lambda}$. Consider the injective homomorphism
$\overline{\Lambda}^{\times}\longrightarrow\mathrm{Aut}_{\mathbb{Z}}(\overline{\Lambda}),\quad x\longmapsto(y\mapsto xy).$
Since we assume that we can compute generators of $\overline{\Lambda}^{\times}$, the problem is reduced to finding a suitable representation for $\mathrm{Aut}_{\mathbb{Z}}(\overline{\Lambda})$. We now describe two ways in which this can be achieved in practice.
  1. (i)
    If $\overline{\Lambda}$ is a free $\mathbb{Z}/n\mathbb{Z}$-algebra of rank $k$ for some $n,k\in\mathbb{Z}_{>0}$, then after choosing a $\mathbb{Z}/n\mathbb{Z}$-basis we have an isomorphism $\mathrm{Aut}_{\mathbb{Z}}(\overline{\Lambda})\cong\mathrm{GL}_{k}(\mathbb{Z}/n\mathbb{Z})$. If we further assume that $n$ is equal to a prime $p$, then this in turn yields a realisation of $\overline{\Lambda}^{\times}$ as a subgroup of $\mathrm{GL}_{k}(\mathbb{F}_{p})$.
  2. (ii)
    Generators of $\mathrm{Aut}_{\mathbb{Z}}(\overline{\Lambda})$ can be computed using results of Shoda [Sho28]. These can be used to determine a faithful permutation representation (see [CH03]) $\mathrm{Aut}_{\mathbb{Z}}(\overline{\Lambda})\longrightarrow\mathfrak{S}_{d}$ for some $d\in\mathbb{Z}_{>0}$, which in turn yields a representation of $\overline{\Lambda}^{\times}$ as a permutation group.
In general, the method of (i) allows computations to be performed more efficiently than the method of (ii), but of course the method of (i) is not always possible. We also note that if $\overline{\Lambda}$ can be written as a direct product of rings $R_{1}\times\cdots\times R_{s}$, then
$\overline{\Lambda}^{\times}=(R_{1}\times\cdots\times R_{s})^{\times}=R_{1}^{\times}\times\cdots\times R_{s}^{\times}$
and so either of the above methods can be applied to each component $R_{i}$ individually, which is more efficient than applying either method to $\overline{\Lambda}$ directly.
12. Improved freeness testing for generalised Swan modules
12.1. Setup
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$ and let $\mathcal{M}$ be a maximal $\mathcal{O}$-order in $A$ containing $\Lambda$. Recall from Definition 8.1 that for $\beta\in\mathcal{M}\cap A^{\times}$, the generalised Swan module $X_{\beta}=X_{\beta,\Lambda,\mathcal{M}}$ is defined to be $\mathcal{M}\beta\cap\Lambda$.
Algorithms 8.9, 9.1 and 10.2 for determining whether $\Lambda$ has SFC each rely on the ability to check whether a given (stably free) $\Lambda$-lattice of the form $X_{\beta}$ is in fact free over $\Lambda$. As explained in the further details of Step (11) of Algorithm 8.9, in principle, this can be performed using [BJ11, Algorithm 3.1]. However, this is only practical when $A$ is of ‘moderate’ dimension. In this section, we shall present a modified freeness test under assumptions of the following paragraph.
Henceforth assume that $A$ admits a decomposition $A=A_{1}\oplus A_{2}$ of $K$-algebras such that $A_{1}$ satisfies the Eichler condition. In applications, we can always assume that $A_{2}\neq 0$ is a direct sum of totally positive quaternion algebras. If $A_{1}=0$ then we need to use a slight modification of [BJ11, Algorithm 3.1] as discussed in the further details of Step (11) of Algorithm 8.9.
Let $e_{1},e_{2}\in A$ be the central idempotents such that $A_{i}=e_{i}A$ for $i=1,2$. Let $C$ denote the centre of $A$ and let $\mathcal{O}_{C}$ denote the (unique) maximal $\mathcal{O}$-order in $C$. Let $\mathfrak{f}$ be any proper full two-sided ideal of $\mathcal{M}$ that is contained in $\Lambda$ and let $\mathfrak{g}=\mathfrak{f}\cap\mathcal{O}_{C}$. Then the central idempotents $e_{1}$ and $e_{2}$ induce decompositions
$\mathcal{M}=\mathcal{M}_{1}\oplus\mathcal{M}_{2},\quad\mathfrak{f}=\mathfrak{f}_{1}\oplus\mathfrak{f}_{2},\quad C=C_{1}\oplus C_{2},\quad\mathcal{O}_{C}=\mathcal{O}_{C_{1}}\oplus\mathcal{O}_{C_{2}},\quad\mathfrak{g}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{2}.$
For $i=1,2$ let $T_{i}$ denote the image of $\mathcal{M}_{i}^{\times}\longrightarrow\left(\mathcal{M}/\mathfrak{f}_{i}\right)^{\times}$ and note that $T:=T_{1}\times T_{2}$ is the image of ×$\mathcal{M}^{\times}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$.
12.2. Criteria for certain ideals to be free
We shall consider left ideals $X$ of $\Lambda$ with the following two properties: $X+\mathfrak{f}=\Lambda$ and there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Note that by Proposition 8.3, the generalised Swan modules $X_{\beta}$ considered in Algorithms 8.9, 9.1 and 10.2 always satisfy both these properties.
Lemma 12.1.
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Write $\beta=(\beta_{1},\beta_{2})\in\mathcal{M}_{1}\oplus\mathcal{M}_{2}$. Then the following are equivalent:
  1. (i)
    $X$ is free over $\Lambda$,
  2. (ii)
    there exists $\overline{\alpha}\in(\Lambda/\mathfrak{f})^{\times}$ such that $\overline{\beta}\overline{\alpha}\in T$,
  3. (iii)
    there exist $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ and $\overline{\alpha}_{1}\in(\mathcal{M}_{1}/\mathfrak{f}_{1})^{\times}$ such that
    $(\overline{\alpha}_{1},\overline{\alpha}_{2})\in(\Lambda/\mathfrak{f})^{\times}\text{ and }\overline{\beta}_{1}\overline{\alpha}_{1}\in T_{1}.$
Proof.
The equivalence of (i) and (ii) follows from Proposition 7.1. Statement (iii) is just an immediate reformulation of (ii). ∎
[BJ11, Algorithm 3.1] essentially uses the the equivalence of conditions (i) and (ii) in Lemma 12.1 to reduce the freeness test to an enumeration over the set $T$ of cardinality $\lvert T\rvert=\lvert T_{1}\rvert\cdot\lvert T_{2}\rvert$. In the following, we show how condition (iii) can be used to replace this by an enumeration over $T_{2}$ and computations related to the reduced norm.
By [BHJ22, Lemma 4.4], there exists a surjective group homomorphism $\overline{\mathrm{nr}}$ that fits into the commutative diagram
${\mathcal{M}_{1}^{\times}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{nr}}$${\mathcal{O}_{C_{1}}^{\times}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\mathcal{M}_{1}/\mathfrak{f}_{1})^{\times}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\overline{\mathrm{nr}}}$${(\mathcal{O}_{C_{1}}/\mathfrak{g}_{1})^{\times},}$ (12.1)
where the vertical maps are induced by the canonical projections. For $i=1,2$, let $\pi_{i}:\Lambda\longrightarrow\mathcal{M}_{i}$ denote the canonical projection and set $\pi_{i}(\Lambda):=\Lambda_{i}$. Hence we can and do also consider $\pi_{i}$ as a map $\pi_{i}\colon:\Lambda\longrightarrow\Lambda_{i}$, depending on context. Note that in fact $\mathfrak{h}_{i}:=\Lambda\cap\mathcal{M}_{i}$ is a two-sided ideal of $\Lambda_{1}\oplus\Lambda_{2}$ contained in $\Lambda$.
Each $\pi_{i}$ induces a group homomorphism
$\overline{\pi}_{i}\colon(\Lambda/\mathfrak{f})^{\times}\longrightarrow(\Lambda_{i}/\mathfrak{f}_{i})^{\times},$ (12.2)
which is surjective by [Bas68, Chapter III, (2.9) Corollary]. Let $\Sigma=\ker(\overline{\pi}_{2})$ and let
$S:=\overline{\mathrm{nr}}(\overline{\pi}_{1}(\Sigma))=\left\{\overline{\mathrm{nr}}(\overline{\gamma}_{1})\mid\exists\overline{\gamma}_{2}\in(\Lambda_{2}/\mathfrak{f}_{2})^{\times}:(\overline{\gamma}_{1},\overline{\gamma}_{2})\in\Sigma\right\}.$ (12.3)
Then $S$ is a subgroup of $\left(\mathcal{O}_{C_{1}}/\mathfrak{g}_{1}\right)^{\times}$ by definition and
$S$ =$=\left\{\overline{\mathrm{nr}}(\overline{\gamma}_{1})\mid\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}\text{ and }(\overline{\gamma}_{1},\overline{1})\in\left(\Lambda/\mathfrak{f}\right)^{\times}\right\},$
=$=\left\{\overline{\mathrm{nr}}(\overline{\gamma}_{1})\mid\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}\text{ and }(\gamma_{1},1)\in\Lambda\right\}$
where the first equality follows from (12.3) and the second follows from the fact that $(\Lambda/\mathfrak{f})^{\times}=(\mathcal{M}/\mathfrak{f})^{\times}\cap(\Lambda/\mathfrak{f})$ (see [Kar92, Chapter I, Theorem 7.9(iii)]). Let $\sigma$ denote the following composition of canonical maps
$\mathcal{O}_{C_{1}}^{\times}\longrightarrow\left(\mathcal{O}_{C_{1}}/\mathfrak{g}_{1}\right)^{\times}\longrightarrow\left(\mathcal{O}_{C_{1}}/\mathfrak{g}_{1}\right)^{\times}/S.$
Lemma 12.2.
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Then $\overline{\beta}\in\left(\mathcal{M}/\mathfrak{f}\right)^{\times}$. Write $\beta=(\beta_{1},\beta_{2})\in\mathcal{M}_{1}\oplus\mathcal{M}_{2}$. Let $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ and assume that $\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ satisfies $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$. Then the following are equivalent:
  1. (i)
    There exists $\overline{\alpha}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $(\overline{\alpha}_{1},\overline{\alpha}_{2})\in(\Lambda/\mathfrak{f})^{\times}\text{ and }\overline{\beta}_{1}\overline{\alpha}_{1}\in T_{1}$,
  2. (ii)
    $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\bmod{S}\in\mathrm{im}(\sigma)$.
Moreover, if these equivalent conditions hold, then $X$ is free over $\Lambda$.
Proof.
We have $\mathcal{M}=\mathcal{M}(X+\mathfrak{f})=\mathcal{M}X+\mathfrak{f}=\mathcal{M}\beta+\mathfrak{f}$, and so $\overline{\beta}\in(\mathcal{M}/\mathfrak{f})^{\times}$.
Assume that (i) holds. Then by definition of $T_{1}$ there exists $u_{1}\in\mathcal{M}_{1}^{\times}$ such that $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\alpha}_{1})=\overline{\mathrm{nr}}(\overline{u}_{1})=\overline{\epsilon}_{1}$ in $\left(\mathcal{O}_{C_{1}}/\mathfrak{g}_{1}\right)^{\times}$, where $\epsilon_{1}:=\mathrm{nr}({u}_{1})$. Hence $\overline{\mathrm{nr}}((\overline{\beta}_{1}\overline{\gamma}_{1})(\overline{\gamma}_{1}^{-1}\overline{\alpha}_{1}))=\overline{\epsilon}_{1}$ and it remains to show that $\overline{\mathrm{nr}}(\overline{\gamma}_{1}^{-1}\overline{\alpha}_{1})\in S$. This in turn is a consequence of
$(\overline{\gamma}_{1}^{-1}\overline{\alpha}_{1},\overline{1})=(\overline{\gamma}_{1}^{-1}\overline{\alpha}_{1},\overline{\alpha}_{2}^{-1}\overline{\alpha}_{2})=(\overline{\gamma}_{1},\overline{\alpha}_{2})^{-1}(\overline{\alpha}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}.$
Assume now that (ii) holds. Then there exists a unit $\epsilon_{1}\in\mathcal{O}_{C_{1}}^{\times}$ and an element $(\overline{\delta}_{1},\overline{1})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ such that $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1}\overline{\delta}_{1})=\overline{\epsilon}_{1}$ in $\left(\mathcal{O}_{C_{1}}/\mathfrak{g}_{1}\right)^{\times}$. By Eichler’s Theorem on units [CR87, (51.34)] there exists $u_{1}\in\mathcal{M}_{1}^{\times}$ such that $\overline{\beta}_{1}\overline{\gamma}_{1}\overline{\delta}_{1}=\overline{u}_{1}$, that is, $\overline{\beta}_{1}\overline{\gamma}_{1}\overline{\delta}_{1}\in T_{1}$. Since $(\overline{\gamma}_{1}\overline{\delta}_{1},\overline{\alpha}_{2})=(\overline{\gamma}_{1},\overline{\alpha}_{2})(\overline{\delta}_{1},\overline{1})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ we may take $\overline{\alpha}_{1}:=\overline{\gamma}_{1}\overline{\delta}_{1}$.
The final assertion follows immediately from Lemma 12.1. ∎
Proposition 12.3.
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Then the following statements hold.
  1. (i)
    $X$ is free over $\Lambda$ if and only if there exist $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ and $\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ and $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\in\mathrm{im}(\sigma)$.
  2. (ii)
    Fix $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$. Suppose there exists $\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ and $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\in\mathrm{im}(\sigma)$. Then for every $\overline{\mu}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $(\overline{\mu}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$, we have $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\mu}_{1})\in\mathrm{im}(\sigma)$.
Proof.
(i) Suppose that $X$ is free over $\Lambda$. Then by Lemma 12.1 there exist $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ and $\overline{\alpha}_{1}\in(\mathcal{M}_{1}/\mathfrak{f}_{1})^{\times}$ such that $(\overline{\alpha}_{1},\overline{\alpha}_{2})\in(\Lambda/\mathfrak{f})^{\times}\text{ and }\overline{\beta}_{1}\overline{\alpha}_{1}\in T_{1}$. Moreover, $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\alpha}_{1})\bmod S\in\mathrm{im}(\sigma)$ by Lemma 12.2, and so we may take $\overline{\gamma}_{1}=\overline{\alpha}_{1}$. The converse holds again by Lemma 12.2.
(ii) The elements $\overline{\beta}_{1}$ and $\overline{\gamma}_{1}$ satisfy Lemma 12.2 (ii). Thus Lemma 12.2 (i) holds, and this is independent of $\overline{\gamma}_{1}$. Thus if $\overline{\mu}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ also satisfies $(\overline{\mu}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$, then Lemma 12.2 (ii) applied to $\overline{\mu}_{1}$ implies that $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\mu}_{1})\bmod S\in\mathrm{im}(\sigma)$. ∎
Lemma 12.4.
Let $\alpha_{2}\in\mathcal{M}_{2}$ with $\overline{\alpha}_{2}\in\left(\mathcal{M}_{2}/\mathfrak{f}\right)^{\times}$. The following are equivalent:
  1. (i)
    There exists an element $\overline{\gamma}_{1}\in(\mathcal{M}_{1}/\mathfrak{f}_{1})^{\times}$ such that $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$,
  2. (ii)
    $\pi_{2}^{-1}(\{\alpha_{2}\})\neq\emptyset$, where $\pi_{2}\colon\Lambda\longrightarrow\mathcal{M}_{2}$ is the canonical projection.
Proof.
It is clear that (i) implies (ii). Suppose that (ii) holds. Then there exists $\alpha_{1}\in\mathcal{M}_{1}$ such that $(\alpha_{1},\alpha_{2})\in\Lambda$. Thus $\alpha_{2}\in\Lambda_{2}$. By assumption $\overline{\alpha}_{2}\in(\mathcal{M}_{2}/\mathfrak{f}_{2})^{\times}$ and hence $\overline{\alpha}_{2}\in\left(\Lambda_{2}/\mathfrak{f}_{2}\right)^{\times}$ since $(\Lambda_{2}/\mathfrak{f}_{2})^{\times}=(\mathcal{M}_{2}/\mathfrak{f}_{2})^{\times}\cap(\Lambda_{2}/\mathfrak{f}_{2})$ (see [Kar92, Chapter I, Theorem 7.9(iii)]). Since the canonical map $\overline{\pi}_{2}\colon(\Lambda/\mathfrak{f})^{\times}\longrightarrow(\Lambda_{2}/\mathfrak{f}_{2})^{\times}$ of (12.2) is surjective, this completes the proof. ∎
Algorithm 12.5.
Suppose that $A_{1}$ is Eichler and that every simple component of $A_{2}$ is a totally definite quaternion algebra. Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Write $\beta=(\beta_{1},\beta_{2})\in\mathcal{M}_{1}\oplus\mathcal{M}_{2}$. If $X$ is free over $\Lambda$ then the following algorithm returns $\mathsf{true}$; otherwise it returns $\mathsf{false}$.
  1. (1)
    Compute $T_{2}$.
  2. (2)
    Compute $S$.
  3. (3)
    Compute $\mathrm{im}(\sigma)$.
  4. (4)
    For each $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ do the following:
    1. (a)
      Compute $\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ or, if no such $\overline{\gamma}_{1}$ exists, continue with the next $\overline{\alpha}_{2}$.
    2. (b)
      If $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\in\mathrm{im}(\sigma)$, return $\mathsf{true}$.
  5. (5)
    Return $\mathsf{false}$.
Proof of correctness.
The correctness of the output follows from Proposition 12.3. ∎
Further details on each step.
Step (1): This can be performed as described in [BJ11, §7].
Step (2): Although the sets $\Sigma$ and $S$ are finite and can be found by enumerating $(\Lambda/\mathfrak{f})^{\times}$ and $(\mathcal{M}_{1}/\mathfrak{f}_{1})^{\times}$, respectively, this is impractical. We present a practical algorithm for computing $\Sigma$ and $S$ in §12.3.
Step (3): This is straightforward given the output of step (2) and standard algorithms from computational algebraic number theory (see [Coh00]).
Step 4(a): For a given $\overline{\alpha}_{2}\in(\mathcal{M}_{2}/\mathfrak{f}_{2})^{\times}$ we can determine whether there exists $\overline{\gamma}_{1}\in(\mathcal{M}_{1}/\mathfrak{f}_{1})^{\times}$ such that $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in(\Lambda/\mathfrak{f})^{\times}$ (and find such an element if it does exist) as follows. The canonical projection $\pi_{2}\colon\Lambda\longrightarrow\mathcal{M}_{2}$ is a ring homomorphism and hence a homomorphism of the underlying finitely generated abelian groups. Thus we can check condition (ii) of Lemma 12.4 and find $\delta_{1}\in\mathcal{M}_{1}$ with $(\delta_{1},\alpha_{2})\in\Lambda$ if it exists (see [Coh00, §4.1]). Note that $(\overline{\delta}_{1},\overline{\alpha}_{2})\in\Lambda/\mathfrak{f}$ maps to $\overline{\alpha}_{2}\in(\Lambda_{2}/\mathfrak{f}_{2})^{\times}$ under the canonical map $\tilde{\pi}_{2}\colon\Lambda/\mathfrak{f}\to\Lambda_{2}/\mathfrak{f}_{2}$, but this preimage of $\overline{\alpha}_{2}$ is not necessarily a unit. Since $\tilde{\pi}_{2}$ is a homomorphism of finite abelian groups, we can determine random elements of $\ker(\tilde{\pi}_{2})$ and thus of $(\overline{\delta}_{1},\overline{\alpha}_{2})+\ker(\tilde{\pi}_{2})=\tilde{\pi}_{2}^{-1}(\{\overline{\alpha}_{2}\})$ until we have found an invertible preimage.
Step 4(b): This can be performed using the commutative diagram (12.1) and standard algorithms for finitely generated abelian groups, as presented for example in [Coh00, §4.1], for example.
Step (5) is trivial. ∎
Remark 12.6.
Compared to [BJ11, Algorithm 3.1] where the size of the enumeration set is essentially the cardinality of the image of the canonical map ×$\mathcal{M}^{\times}\longrightarrow\left(\mathcal{M}/\mathfrak{f}\right)^{\times}$, that is, $\lvert T_{1}\rvert\cdot\lvert T_{2}\rvert$, the enumeration set in Algorithm 12.12 is of size $\lvert T_{2}\rvert$, and hence only depends on the component $A_{2}$. Note that $T_{2}$ can be computed as explained in [BJ11, §4.5] (see also [KV10, Remark 7.5]), and this is much faster than computing the image of $\mathcal{M}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$ for $\mathcal{M}$ a maximal $\mathcal{O}$-order in an arbitrary finite-dimensional semisimple $K$-algebra.
12.3. Computation of $S$
Recall the definition of $S$ given in (12.3). Since we require only generators of $S$, it will be enough to consider the computation of (generators of) $\Sigma=\ker(\overline{\pi}_{2})$. Recall that $\mathfrak{h}_{1}=\Lambda\cap\mathcal{M}_{1}$ and consider the two-sided ideal $\tilde{\mathfrak{f}}:=\mathfrak{h}_{1}+\mathfrak{f}=\mathfrak{h}_{1}\oplus\mathfrak{f}_{2}$. Let
$\overline{\pi}\colon(\Lambda/\mathfrak{f})^{\times}\longrightarrow(\Lambda/\tilde{\mathfrak{f}})^{\times}$
denote the canonical map, which is surjective by [Bas68, Chapter III, (2.9) Corollary].
Lemma 12.7.
We have $\Sigma=\ker(\overline{\pi})$.
Proof.
Since the composition $\Lambda\stackrel{{\scriptstyle\pi_{2}}}{{\longrightarrow}}\Lambda_{2}\longrightarrow\Lambda_{2}/\mathfrak{f}_{2}$ is surjective with kernel $\tilde{\mathfrak{f}}$, we obtain an isomorphism $\overline{\tau}\colon(\Lambda/\tilde{\mathfrak{f}})^{\times}\longrightarrow(\Lambda_{2}/\mathfrak{f}_{2})^{\times}$ making the diagram
${(\Lambda/\tilde{\mathfrak{f}})^{\times}}$${(\Lambda/\mathfrak{f})^{\times}}$${(\Lambda_{2}/\mathfrak{f}_{2})^{\times}}$$\scriptstyle{\overline{\tau}}$$\scriptstyle{\overline{\pi}}$$\scriptstyle{\overline{\pi}_{2}}$
commute. The claim now follows from $\Sigma=\ker(\overline{\pi}_{2})$. ∎
Lemma 12.8.
There exist coprime primary ideals $\{\mathfrak{q}_{1},\dotsc,\mathfrak{q}_{r+s}\}$ and $\{\tilde{\mathfrak{q}}_{1},\dotsc,\tilde{\mathfrak{q}}_{r}\}$ of $\Lambda\cap C$ such that the following hold:
  1. (i)
    we have $\mathfrak{q}_{i}\subseteq\tilde{\mathfrak{q}}_{i}$ for $1\leq i\leq r$,
  2. (ii)
    we have
    $\Lambda/\mathfrak{f}\cong\prod_{i=1}^{r+s}\Lambda/(\mathfrak{q}_{i}\Lambda+\mathfrak{f})\quad\text{and}\quad\Lambda/\tilde{\mathfrak{f}}\cong\prod_{i=1}^{r}\Lambda/(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}}),$
  3. (iii)
    we have
    $\Sigma\cong\prod_{i=1}^{r}\ker(\tau_{i})\times\prod_{i=1}^{s}(\Lambda/(\mathfrak{q}_{r+i}\Lambda+\mathfrak{f}))^{\times},$
    where $\tau_{i}\colon(\Lambda/(\mathfrak{q}_{i}\Lambda+\mathfrak{f}))^{\times}\longrightarrow(\Lambda/(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}}))^{\times}$ are the canonical surjections.
Proof.
Consider $\mathfrak{g}=\mathfrak{f}\cap C$ and $\tilde{\mathfrak{g}}=\mathfrak{\tilde{f}}\cap C$, which are full $(\Lambda\cap C)$-ideals with $\mathfrak{g}\subseteq\tilde{\mathfrak{g}}$. Let $\mathfrak{p}_{1},\dotsc,\mathfrak{p}_{r}$ and $\mathfrak{p}_{1},\dotsc,\mathfrak{p}_{r+s}$ be the maximal ideals of $\Lambda\cap C$ containing $\tilde{\mathfrak{g}}$ and $\mathfrak{g}$ respectively. Let now $\mathfrak{g}=\prod_{i=1}^{r+s}\mathfrak{q_{i}}$ be a primary decomposition, where $\mathfrak{q}_{i}$ is $\mathfrak{p}_{i}$-primary for $1\leq i\leq r+s$. Since the $\mathfrak{p}_{i}$ are maximal, it is clear that for each $1\leq i\leq r$ also $\mathfrak{q}_{i}+\mathfrak{\tilde{g}}$ is $\mathfrak{p}_{i}$-primary. Thus $\tilde{\mathfrak{g}}=\mathfrak{g}+\tilde{\mathfrak{g}}=\prod_{i=1}^{r+s}\mathfrak{q}_{i}+\tilde{\mathfrak{g}}=\prod_{i=1}^{r}\tilde{\mathfrak{q}}_{i}$ is a primary decomposition of $\tilde{\mathfrak{g}}$, where $\tilde{\mathfrak{q_{i}}}:=\mathfrak{q}_{i}+\tilde{\mathfrak{g}}$ is $\mathfrak{p}_{i}$-primary for $1\leq i\leq r$. This immediately shows (i). For (ii), note that [BB06, 3.5 Lemma] implies that
$\mathfrak{f}=\bigcap_{i=1}^{r+s}(\mathfrak{q}_{i}\Lambda+\mathfrak{f})\quad\text{ and }\quad\tilde{\mathfrak{f}}=\bigcap_{i=1}^{r}(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}}),$
thus the claim follows from the Chinese remainder theorem. Finally, (iii) follows from (ii) together with Lemma 12.7. ∎
Since we know how to compute generators for each of the groups $\left(\Lambda/(\mathfrak{q}_{r+i}\Lambda+\mathfrak{f})\right)^{\times}$ by the results of §11.2, it remains to consider the computation of $\ker(\tau_{i})$ for $i=1,\ldots,r$. Each element in $\ker(\tau_{i})$ is the image in $(\Lambda/(\mathfrak{q}_{i}\Lambda+\mathfrak{f}))^{\times}$ of an element of the form $1+\lambda$ with $\lambda\in(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}})$. Note, however, that not every element in $1+(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}})$ is actually a unit mod $(\mathfrak{q}_{i}\Lambda+\mathfrak{f})$.
Lemma 12.9.
Let $\mathfrak{p}$ be a maximal ideal of $\Lambda\cap C$, let $\mathfrak{q}$ and $\tilde{\mathfrak{q}}$ be two $\mathfrak{p}$-primary ideals of $\Lambda\cap C$ with ${\mathfrak{q}}\subseteq\tilde{\mathfrak{q}}$ and let
$\tau\colon(\Lambda/(\mathfrak{q}\Lambda+\mathfrak{f}))^{\times}\longrightarrow(\Lambda/(\tilde{\mathfrak{q}}\Lambda+\tilde{\mathfrak{f}}))^{\times}$
the canonical projection. Furthermore set
$\tilde{\mathfrak{Q}}:=\tilde{\mathfrak{q}}\Lambda+\tilde{\mathfrak{f}},\quad\mathfrak{Q}_{1}:=\tilde{\mathfrak{q}}\Lambda+\mathfrak{f},\quad\mathfrak{Q}:=\mathfrak{q}\Lambda+\mathfrak{f}.$
Then the following hold:
  1. (i)
    for every $\alpha\in\Lambda$ we have
    $\alpha\text{ is invertible mod }\mathfrak{Q}\iff\alpha\text{ is invertible mod}\mathfrak{Q}_{1}.$
  2. (ii)
    Identifying $(1+\tilde{\mathfrak{Q}})$ and $(1+\mathfrak{Q}_{1})$ with their images in $\Lambda/\mathfrak{Q}$ we have $\ker(\tau)=(1+\tilde{\mathfrak{Q}})\cap\left(\Lambda/\mathfrak{Q}\right)^{\times}$ and the sequence
    $0\longrightarrow\frac{1+\mathfrak{Q}_{1}}{1+\mathfrak{Q}}\longrightarrow\ker(\tau)\longrightarrow\frac{(1+\tilde{\mathfrak{Q}})\cap\left(\Lambda/\mathfrak{Q}\right)^{\times}}{1+\mathfrak{Q}_{1}}\longrightarrow 0$
    is exact.
Proof.
(i) Since both $\tilde{\mathfrak{q}}$ and $\mathfrak{q}$ are $\mathfrak{p}$-primary, there exists $m>0$ such that $\tilde{\mathfrak{q}}^{m}\subseteq\mathfrak{q}$. Hence $\mathfrak{Q}_{1}^{m}\subseteq\mathfrak{Q}$ and every element of $\mathfrak{Q}_{1}/\mathfrak{Q}$ is nilpotent in $\Lambda/\mathfrak{Q}$. If $\lambda$ is invertible mod $\mathfrak{Q}_{1}$, then there exist $\mu\in\Lambda$ and $\xi\in\mathfrak{Q}_{1}$ such that $\lambda\mu=1-\xi$. Thus in $\Lambda/\mathfrak{Q}$ we have $\overline{\lambda}\overline{\mu}\cdot\sum_{n=0}^{\infty}\overline{\xi}^{n}=(\overline{1}-\overline{\xi})\cdot\sum_{n=0}^{\infty}\overline{\xi}^{n}=\overline{1}$. The converse is clear. (ii) now follows easily from (i). ∎
Remark 12.10.
In conjunction with Lemma 12.8, Lemma 12.9 yields the following method for computing generators of $\Sigma$ (and hence of $S$). Primary decompositions in $\Lambda\cap C$ can be determined using Corollary 11.3 and thus a decomposition
$\Sigma\cong\prod_{i=1}^{r}\ker(\tau_{i})\times\prod_{i=1}^{s}(\Lambda/(\mathfrak{q}_{r+i}\Lambda+\mathfrak{f}))^{\times},$
with suitable primary ideals $\mathfrak{q}_{i},\tilde{\mathfrak{q}}_{i}$ as in Lemma 12.8 (iii) can be obtained. For the groups $(\Lambda/(\mathfrak{q}_{r+i}\Lambda+\mathfrak{f}))^{\times}$, generators can be obtained by Corollary 11.4 and thus it suffices to determine each $\ker(\tau_{i})$. For $1\leq i\leq r$ we set $\tau:=\tau_{i}$ and fix the notation as in Lemma 12.9. Since
$1\longrightarrow\frac{1+\mathfrak{Q}_{1}}{1+\mathfrak{Q}}\longrightarrow\ker(\tau)\longrightarrow\frac{(1+\tilde{\mathfrak{Q}})\cap\left(\Lambda/\mathfrak{Q}\right)^{\times}}{1+\mathfrak{Q}_{1}}\longrightarrow 1$ (12.4)
is exact, it suffices to find generators for the first and third group. Generators for $\frac{1+\mathfrak{Q}_{1}}{1+\mathfrak{Q}}$ can be computed as in [BB06, §3.7]. For the computation of generators of the non-trivial group on the right hand side of (12.4), the elements can be listed as follows: determine the set of representatives of $\tilde{\mathfrak{Q}}/\mathfrak{Q}_{1}$ and check for each such representative $\alpha$ whether $1+\alpha$ it is a unit mod $\mathfrak{Q}$.
12.4. A probabilistic algorithm
In the previous section, the computation of $\Sigma$ respectively $S$ was reduced to the problem of enumerating all elements of quotients of the form $\tilde{\mathfrak{Q}}/\mathfrak{Q}_{1}$ for two-sided $\Lambda$-ideals $\tilde{\mathfrak{Q}}$ and $\mathfrak{Q}_{1}$. Since the order of $\tilde{\mathfrak{Q}}/\mathfrak{Q}_{1}$ can be quite large, we present a probabilistic approach, which avoids enumeration. We assume that $S^{\prime}\subseteq S$ is a subgroup and consider the following composition, which we denote by $\sigma^{\prime}$:
$\mathcal{O}_{C_{1}}^{\times}\longrightarrow\left(\mathcal{O}_{C_{1}}/\mathfrak{g}_{1}\right)^{\times}\longrightarrow{\left(\mathcal{O}_{C_{1}}/\mathfrak{g}_{1}\right)^{\times}}/S^{\prime}.$
Replacing $S$ by $S^{\prime}$ in Lemma 12.2, we obtain the following weaker statement.
Lemma 12.11.
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Write $\beta=(\beta_{1},\beta_{2})\in\mathcal{M}_{1}\oplus\mathcal{M}_{2}$. Let $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ and assume that $\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ satisfies $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$. If $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\in\mathrm{im}(\sigma^{\prime})$, then $X$ is free over $\Lambda$.
This yields the following probabilistic variant of Algorithm 12.12.
Algorithm 12.12.
Suppose that $A_{1}$ is Eichler and that every simple component of $A_{2}$ is a totally definite quaternion algebra. Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. If the following algorithm returns $\mathsf{true}$, then $X$ is free over $\Lambda$; if returns $\mathsf{inconclusive}$, then nothing can be said about whether $X$ is free over $\Lambda$ or not.
  1. (1)
    Compute $T_{2}$.
  2. (2)
    Compute $S^{\prime}$.
  3. (3)
    Compute $\mathrm{im}(\sigma^{\prime})$.
  4. (4)
    For each $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ do the following:
    1. (a)
      Compute $\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ or, if no such $\overline{\gamma}_{1}$ exists, continue with the next $\overline{\alpha}_{2}$.
    2. (b)
      If $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\in\mathrm{im}(\sigma^{\prime})$, return $\mathsf{true}$.
  5. (5)
    Return $\mathsf{inconclusive}$.
Proof of correctness.
The correctness of the output follows from Lemma 12.11. ∎
Further details on each step.
Apart from Step (2), all steps are the essentially the same as in Algorithm 12.12. Step (2) can be performed as follows. Recall that $S=\overline{\mathrm{nr}}(\overline{\pi}_{1}(\Sigma))$. Thus in order to find a subgroup $S^{\prime}$ of $S$ it is sufficient to find a subgroup of
$\Sigma=\ker((\Lambda/\mathfrak{f})^{\times}\longrightarrow(\Lambda_{2}/\mathfrak{f}_{2})^{\times})).$
Random elements of $\Sigma$ can be found by sampling preimages of $\overline{1}\in\Lambda_{2}/\mathfrak{f}_{2}$ under the canonical map $\Lambda/\mathfrak{f}\longrightarrow\Lambda_{2}/\mathfrak{f}_{2}$ (a homomorphism of finitely generated abelian groups) and discarding the non-invertible elements. In this way we obtain a sequence of elements $(\overline{\alpha}_{i})_{i\in\mathbb{Z}_{\geq 1}}$ in $\Sigma$ and determine $S^{\prime}:=\langle\overline{\mathrm{nr}}(\bar{\pi}_{1}(\bar{\alpha}_{i}))\mid i\leq B\rangle$, where $B\in\mathbb{Z}_{\geq 1}$ is either a fixed predetermined bound or chosen dynamically until $S^{\prime}$ stabilises. ∎
13. Further results on SFC for integral group rings
Theorem 13.1.
Let $G$ be one of the following finite groups:
$\tilde{T}\times C_{2},\,\tilde{T}\times Q_{12},\,\tilde{T}\times Q_{20},\,Q_{8}\rtimes\tilde{T},\,G_{(96,66)},\,G_{(192,183)},\,G_{(480,962)}.$
Then $\mathbb{Z}[G]$ has SFC.
Proof.
The result for $\tilde{T}\times C_{2}$ is Theorem 6.6. Note that $\tilde{T}\times Q_{12}$ and $\tilde{T}\times Q_{20}$ both have quotients isomorphic to $\tilde{T}\times C_{2}$, so Theorem 6.6 can be recovered from Lemma 4.1 and the result for either $\tilde{T}\times Q_{12}$ or $\tilde{T}\times Q_{20}$.
For $G=G_{(96,66)}$, $\tilde{T}\times Q_{12}$ or $\tilde{T}\times Q_{20}$, there exists a normal subgroup $N$ of order $2$, such that $H:=G/N$ is isomorphic to $\mathrm{GL}_{2}(\mathbb{F}_{3})$, $G_{(144,127)}$, or $G_{(280,108)}$, respectively. In each case, $\mathbb{Q}[H]$ satisfies the Eichler condition and hence so does $\mathbb{Q}[H\times C_{2}]$. Thus $\mathbb{Z}[H\times C_{2}]$ has LFC by Theorem 3.4. Let $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$. The cardinalities of the locally free class groups of $\mathbb{Z}[H]$, $\mathbb{Z}[G]$, $\mathbb{Z}[H\times C_{2}]$ and $\Gamma$ can be computed using the the algorithm of [BB06, §3], and thus it can be verified that Proposition 6.10 can be applied to show that $\mathbb{Z}[G]$ has SFC if only if $\Gamma$ has SFC.
For $G=Q_{8}\rtimes\tilde{T}\cong G_{(192,1022)}$, there is a normal subgroup $N$ of order $2$ such that $H:=G/N$ is isomorphic to $G_{(96,204)}$. It can easily be checked that $\mathbb{Q}[H]$ satisfies the Eichler condition. Let $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$. Unfortunately, Proposition 6.10 cannot be applied in this situation because the required equality (6.3) does not hold. However, $\mathbb{Z}[G]$ can be shown to have SFC if and only if $\Gamma$ has SFC by an application of Algorithm 10.3 with $\Lambda_{1}=\mathbb{Z}[H]$ and $\Lambda_{2}=\Gamma$. Note that $\overline{\Lambda}=\mathbb{F}_{2}[H]$, and so an explicit embedding of $\overline{\Lambda}^{\times}$ into $\mathrm{GL}_{96}(\mathbb{F}_{2})$ can be determined as explained in Remark 6.3 (see also §11.3). Crucially, this allows computations in $\overline{\Lambda}^{\times}$ to be performed reasonably efficiently.
Therefore, for $G=G_{(96,66)}$, $G_{(192,1022)}$, $\tilde{T}\times Q_{12}$, or $\tilde{T}\times Q_{20}$, we have shown that $\mathbb{Z}[G]$ has SFC if and only if $\Gamma$ has SFC. This is highly advantageous from a computational point of view, since $\dim_{\mathbb{Q}}\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma=\frac{1}{2}\dim_{\mathbb{Q}}\mathbb{Q}[G]=\frac{1}{2}|G|$. To verify that $\Gamma$ has SFC, we again apply Algorithm 10.3, this time with the fibre product induced by the decomposition $\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma=B_{1}\oplus B_{2}$, with $B_{1}$ the largest component of $\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma$ satisfying the Eichler condition (equivalently, with $B_{2}$ the largest component that is the direct sum of totally definite quaternion algebras). Computations in $\overline{\Lambda}^{\times}$ are performed by first computing an explicit embedding into the symmetric group $\mathfrak{S}_{d}$ for some $d\in\mathbb{Z}_{>0}$ (see §11.3). This is less efficient than working with an embedding into $\mathrm{GL}_{n}(\mathbb{F}_{p})$ for appropriate choices of $n$ and $p$, but such an embedding does not exist in general.
For $G=G_{(192,183)}$ or $G_{(480,962)}$, there exists a normal subgroup $N$ of such that $H:=G/N\cong G_{(96,66)}$ and so $\mathbb{Z}[H]$ has SFC by the result for $G_{(96,66)}$. Let $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$. It can easily be checked that $\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma$ satisfies the Eichler condition and thus $\Gamma$ has LFC by Theorem 3.4. Then $\mathbb{Z}[G]$ can be shown to have SFC by an application of Algorithm 10.3 with $\Lambda_{1}=\Gamma$ and $\Lambda_{2}=\mathbb{Z}[H]$. Note that $\overline{\Lambda}=\mathbb{F}_{p}[H]$ for $p=2$ or $5$, which allows computations in $\overline{\Lambda}^{\times}$ to be performed reasonably efficiently, as explained above. ∎
Theorem 13.2.
Let $G$ be a finite group with $|G|\leq 1023$. Let
$\Sigma=\{3595,3599,3704,3716,3719,3721,3722,4179,4487,4490,5069,8602,8604,8614\}.$
Suppose that $G$ is not isomorphic to any of the following groups:
  1. (a)
    $G_{(384,m)}$ for $m\in\{572,577,579,5867,18129,18229\}$,
  2. (b)
    $G_{(576,m)}$ for $m\in\{1447,2006\}$,
  3. (c)
    $G_{(768,m+1080000)}$ for $n\in\Sigma$,
  4. (d)
    $G_{(864,m)}$ for $m\in\{717,731,732,3021\}$, or
  5. (e)
    $G_{(960,805)}$.
Then $\mathbb{Z}[G]$ has SFC if and only if $G$ has no quotient isomorphic to
  1. (i)
    $Q_{4n}\times C_{2}$ for $2\leq n\leq 5$,
  2. (ii)
    $\tilde{T}\times C_{2}^{2}$, $\tilde{O}\times C_{2}$, $\tilde{I}\times C_{2}$,
  3. (iii)
    $G_{(32,14)}$, $G_{(36,7)}$, $G_{(64,14)}$, $G_{(100,7)}$, or
  4. (iv)
    $Q_{4n}$ for $6\leq n\leq 255$.
Proof.
We use the GAP [GAP24] small groups database [BEO02], which in particular, contains all finite groups $G$ with $|G|\leq 1023$. For each such $G$, we first check whether $\mathbb{Z}[G]$ has SFC by Theorems 4.5, 4.10, 4.11, and 13.1, as well as Corollary 6.9. We then check whether $G$ has a quotient isomorphic to any of the groups listed in (i), (ii), (iii), or (iv), in which case $\mathbb{Z}[G]$ fails SFC by Corollary 9.7. These checks can easily be implemented as a program in GAP [GAP24] or Magma [BCP97]. For $G_{(480,256)}\cong C_{5}\times G_{(96,66)}$, $G_{(576,7442)}\cong C_{3}\times(Q_{8}\rtimes\tilde{T})$, and $G_{(960,9669)}\cong C_{5}\times(Q_{8}\rtimes\tilde{T})$, combination of Theorems 4.6 and 13.1 shows that $\mathbb{Z}[G]$ has SFC. The remaining groups consist of all those listed in (a)–(e). ∎
Corollary 13.3.
Let $G$ be a finite group with $|G|\leq 383$. Then $\mathbb{Z}[G]$ has SFC if and only if $G$ has no quotient of the form
  1. (i)
    $Q_{4n}\times C_{2}$ for $2\leq n\leq 5$,
  2. (ii)
    $\tilde{T}\times C_{2}^{2}$, $\tilde{O}\times C_{2}$, $\tilde{I}\times C_{2}$,
  3. (iii)
    $G_{(32,14)}$, $G_{(36,7)}$, $G_{(64,14)}$, $G_{(100,7)}$, or
  4. (iv)
    $Q_{4n}$ for $6\leq n\leq 95$.